### v1

# MHRC-Bench: A Multilingual Hardware Repository-Level Code Completion benchmark

Abstract
Large language models (LLMs) have achieved strong performance on code completion tasks in general-purpose programming languages. However, existing repository-level code completion benchmarks focus almost exclusively on software code and largely overlook hardware description languages. In this work, we present MHRC-Bench, consisting of MHRC-Bench-Train and MHRC-Bench-Eval, the first benchmark designed for multilingual hardware code completion at the repository level. Our benchmark targets completion tasks and covers three major hardware design coding styles. Each completion target is annotated with code-structure-level and hardware-oriented semantic labels derived from concrete syntax tree analysis. We conduct a comprehensive evaluation of models on MHRC-Bench-Eval. Comprehensive evaluation results and analysis demonstrate the effectiveness of MHRC-Bench.
MHRC-Bench: A Multilingual Hardware Repository-Level Code Completion benchmark
Qingyun Zou1, Jiahao Cui1, Nuo Chen1, Bingsheng He1, Weng-Fai Wong1 1National University of Singapore qingyunzou@u.nus.edu
Figure 1: Average EM and ES comparison of different LLMs on the MHRC-Bench-Eval benchmark.
Figure 2: Overview of our hardware code completion benchmark across multiple hardware description paradigms. We present three representative examples from generator-based hardware construction languages (e.g., Chisel), high-level synthesis (HLS) languages (e.g., Xilinx HLS), and Register-Transfer Level (RTL) design languages (SystemVerilog). For each example, both in-file and cross-file contexts are provided, with the completion target explicitly marked. Models are required to predict the correct completion given the provided contexts, where “<TARGET>” indicates the completion position. More examples are introduced in Appendix A.10
1 Introduction
Recent advancements in large language models (LLMs) have led to significant progress in code-related tasks such as code understanding, generation, and completion (Liu et al., 2023; Nam et al., 2024; Liu et al., 2024a; Pan et al., 2024; Raihan et al., 2025; Yang et al., 2024).
These models, trained on vast datasets encompassing both natural language and source code, have demonstrated remarkable proficiency in modeling programming languages as structured sequences and predicting logical code continuations (Liang et al., 2024; Li et al., 2025; Guo et al., 2024; Hui et al., 2024; Liu et al., 2024b; Abedu et al., 2025; Hua et al., 2025). Consequently, LLM-based code completion has emerged as a valuable tool in real-world software development environments.
More recently, research efforts have shifted toward repository-level code completion, where models are required to leverage cross-file information such as imports, symbol definitions, and project structure. Wang et al. (2024) train a dedicated retriever to identify and retrieve relevant code fragments that can be incorporated into LLM-based code completion. In a complementary direction, Deng et al. (2025) equip LLMs with external tool APIs to more effectively utilize repository-level context during the code completion process. While recent work has explored Verilog code completion (Wu et al., 2025b; Allam and Shalan, 2024), existing benchmarks remain narrowly focused on a single hardware language. Inspired by Liu et al. (2025b), we propose MHRC-Bench, the first multilingual repository-level benchmark for hardware code completion, consisting of MHRC-Bench-Train and MHRC-Bench-Eval.
Using the MHRC-Bench-Train dataset to fine-tune code LLMs, we observe that billion-scale fine-tuned models such as Qwen2.5-Coder-7B achieve substantially higher EM and ES scores than their pretrained counterparts, whose EM is nearly zero on hardware code completion tasks. After fine-tuning, these models also outperform larger general-purpose models, including GPT-5, Grok-4, Gemini 2.5 Pro, and DeepSeek V3.2, as shown in Figure 1. This demonstrates that MHRC-Bench-Train effectively enhances the hardware code completion capabilities of smaller-scale code LLMs.
As shown in Figure 2, MHRC-Bench covers three main categories of hardware programming languages. To enable fine-grained analysis of LLM performance on hardware code completion, MHRC-Bench-Eval introduces two annotation levels: code-structure and hardware semantics. At the code-structural level, source files are parsed into concrete syntax trees (CSTs) using tree-sitter, and completion targets are assigned to depth-based buckets according to the enclosing CST node, capturing structural complexity. At the hardware-semantic level, code is categorized into nine subcategories, enabling detailed evaluation across diverse hardware semantics.
The benchmark and related code are publicly available. The contributions are summarized as follows:
  • We introduce MHRC-Bench-Eval, the first large-scale multilingual repository-level hardware code completion benchmark, covering three major hardware programming paradigms across the hardware design flow. The benchmark provides both code-structure-level and hardware-semantic-level annotations derived from syntax tree analysis.
  • We construct MHRC-Bench-Train, a training dataset designed to enhance repository-level hardware code completion. Fine-tuning on MHRC-Bench-Train leads to substantial performance gains, even enabling smaller open-weight models to outperform strong commercial baselines.
  • We conduct comprehensive evaluations across a wide range of models, including commercial LLMs and open-source code LLMs, as well as state-of-the-art retrieval-based methods. Our analysis reveals key performance trends and insights into LLM behavior in hardware code completion.
2 Related Works and Background
2.1 Repository-level Code Completion.
Recent repository-level code completion methods (Deng et al., 2025; Li et al., 2025; Liu et al., 2024b; Liang et al., 2024; Liu et al., 2025d; Wu et al., 2025b; Wang et al., 2024; Zhang et al., 2025) aim to retrieve related code snippets across files within a repository, where existing datasets mainly focus on software engineering task. For example, M2RC-Eval (Liu et al., 2025b) supports Java, Python and so on these related software codes completion.
2.2 LLM for EDA.
Recent advancements in applying large language models (LLMs) to electronic design automation (EDA) tasks, such as RTL/HLS code generation and hardware verification, have shown significant promise. Notable studies in NLP venues include SynC-LLM (Liu et al., 2025c), which uses diffusion models and LLM prompting to generate realistic digital circuits, improving downstream EDA model performance. Wu et al. (2025a) introduces a multi-agent LLM system that enhances the reliability of multi-step EDA tool flows. DeepRTL2 (Liu et al., 2025e) showcases LLMs’ versatility in RTL generation and embedding-based tasks, setting new benchmarks. Lastly, AssertionBench (Pulavarthi et al., 2025) presents a benchmark for Verilog designs and assertion generation, highlighting progress and areas for improvement. These studies demonstrate the growing impact of LLMs in automating hardware design. However, these works only focus on kernel-level code rather than repository-level codes and only concentrate on one hardware design language.
3 Dataset
3.1 Data Collection
We collect hardware design repositories from GitHub under permissive open-source licenses (e.g., MIT), created after 2000 and with more than five GitHub stars. Repositories with copyleft or restrictive licenses (e.g., GPL) are excluded. To avoid data leakage, we remove files located in common vendor or dependency directories. Also, we ensure dataset uniqueness by removing exact duplicate repositories. The resulting corpus contains 47,175 source files written in Chisel, SystemVerilog, Verilog, HLS C/C++, and VHDL, drawn from 584 distinct repositories
3.2 Dataset Comparison
Language scope. As shown in Table 1 and Figure 2, we select Verilog/SystemVerilog, VHDL, High Level Synthesis, and Chisel to construct our dataset, as they cover three complementary categories:
  • Register Transfer Level (RTL) based hardware description languages. They provide cycle-accurate and explicitly timed hardware descriptions including Verilog/SystemVerilog and VHDL. SystemVerilog and Verilog share similar grammatical structures, whereas VHDL differs significantly. Therefore, we group Verilog and SystemVerilog together (V/SV);
  • High-Level Synthesis (HLS) languages. They describe hardware behavior using C/C++ like abstractions augmented with synthesis directives such as Xilinx HLS C/C++ and SystemC;
  • Generator-based hardware construction languages. They enable parameterized and programmable generation of RTL designs such as Chisel.
Compared with existing hardware repository-level benchmark-RTL-Repo Allam and Shalan (2024), MHRC-Bench enables a comprehensive evaluation of LLMs across different abstraction levels, programming styles, and domain-specific constraints in hardware-oriented code completion.
Table 1: Comparison with repository-level hardware code datasets. Lang indicates languages, FG indicates fine-grained annotation.
Benchmark Lang FG Target Samples
RTL-Repo RTL × Single-line 4098
Ours RTL, HLS, Generator Single & Multi-line 47,175
3.3 Dataset Quality Control.
The selection of completion targets is critical to the quality and reliability of a code completion benchmark. Accordingly, we adopt the following rules to identify and select target code locations.
Figure 3: Illustration of fine-grained completion target annotation in our benchmark. Completion targets are identified by aligning source code positions with Tree-sitter CST nodes and labeled with structure depth and hardware-oriented semantic categories.
Syntactically complete. Prior work Ding et al. (2023) often selects random character spans as completion targets, which may break syntactic integrity. In contrast, recent studies (Liu et al., 2025b) show that developers expect LLMs to generate syntactically complete code units (e.g., full lines or blocks). Following this insight, MHRC-Bench-Eval uses tree-sitter [^1] https://github.com/tree-sitter/tree-sitter to parse each file into a concrete syntax tree and randomly selects a syntactic node as the completion target, using its exact source span as the ground-truth output.
Data Independence and Inference difficulty. To ensure data independence and inference difficulty: (1) The repositories and files used for the training, validation, and test sets are mutually exclusive, with no overlap. (2) We ensure that 40% of completion targets are between 2 and 5 lines inclusive. (3) The target code should not exclusively include whitespace, debug message and comments. (4) For each file, we select exactly one target position.
3.4 Dataset Statistics
Training, Validation, and Test Dataset Splits. To prevent data leakage, the training, validation, and test sets are constructed from disjoint repositories, ensuring that no repository appears in more than one split. This repository-level separation avoids sharing context across splits and enables a fair evaluation of generalization. The sizes of the training, validation, and test sets are shown in Table 2.
Cross-file Dependencies. A key distinction between repository-level code generation and single-file code generation lies in the presence of cross-file dependencies. To quantify this structural complexity, we measure the average number of cross-file dependencies per task across different hardware description languages, as shown in Table 2. The results indicate that the average dependency count is consistently greater than one, demonstrating that resolving references across multiple files is common and essential. This highlights the importance of modeling cross-file context for realistic hardware code completion tasks. The procedure for computing cross-file dependencies is described in Appendix A.1.
Table 2: Dataset statistics of MHRC-Bench. Val means validation. Cross-file means cross-file dependencies. Every column lists its numbers of completion targets.
Language Repos Train Val Test Cross-file
Chisel 48 2,573 119 319 3.12
Verilog/SV 247 25,586 107 492 4.02
VHDL 239 13,039 104 491 4.34
HLS 50 3,766 123 456 5.85
Table 3: Hardware-oriented semantic categories with language-specific representative constructs.
Category SystemVerilog / Verilog VHDL Chisel HLS C/C++
Design Structure “module”, “package” “entity”, “architecture” “Module”, “Bundle” “Top Function”, “Kernel”
Declaration and Definition “wire”, “reg” “signal”, “generic” “IO”, “Wire” “Function Signature”
Storage Block “reg”, “always_ff” “clocked signal”, “process” “Reg”, “Mem” “Array”, “hls::stream”
Computation Block “Arithmetic Op”, “Bit Slice” “Arithmetic Op”, “Type Conversion” “Arithmetic Op”, “Mux” “Arithmetic Expression”, “Array Access”
Control Flow Block “always”, “if” “process”, “if” “when”, “switch” “for loop”, “if”
Interface “Port List”, “Ready/Valid” “Port Map”, “Component Instantiation” “Decoupled”, “Valid” “AXI Interface”, “Stream Interface”
Property and Assertion Specification “assert”, “property” “assert”, “PSL” “assert” “assert”, “check”
Testbench Stimulus and Environment “initial”, “always” “process”, “wait” “poke”, “step” “testbench loop”
Monitoring and Checking Logic “display”, “error” “report”, “assert” “expect” “golden check”
3.5 Dataset Categories
To better analyze LLM’s performance in a fine-grained manner, we further provide two types of fine-grained annotations include code structural level and hardware semantic level) for each completion target as shown in Figure 3.
Code structural level. We first parse the source code into a concrete syntax tree (CST). From the CST, we select syntactically complete nodes as candidate target code regions for completion. For each selected node, we get its depth in the tree, which reflects its structural granularity within the program. Then we divide each tree into M buckets based on the depth degree of the concrete syntax tree. we set M to 5 in MHRC-Bench-Eval.
Hardware semantic level. We categorize the target code regions according to the hardware design and verification flow in Table 3, which broadly consists of design coding and testbench-based verification. Unlike software programs, where execution is inherently sequential, hardware code describes a parallel architecture or circuit with explicit structural, temporal, and interface semantics. This fundamental difference motivates a hardware-specific categorization of completion targets.
For design coding, we divide the target code into the following categories: Design Structure, Declaration and Definition, Storage Block, Computation Block, Control Flow Block, and Interface. These categories capture the key aspects of hardware design, ranging from structural hierarchy and state storage to data-path computation, control logic, communication interfaces.
In addition, we explicitly model verification code, which plays a critical role in hardware development but is largely absent in software benchmarks. We further decompose the verification code into several sub-types, including Property and Assertion Specification, Testbench Stimulus and Environment Modeling, and Monitoring and Checking Logic. These sub-types collectively describe how correctness properties are specified and how its behavior is observed and validated during simulation or formal verification. The distribution of hardware semantic categories is presented in Appendix A.2.
4 Experiments
4.1 Evaluation Models and Metrics
Code model. We evaluate Code LLMs of different scale. The evaluated models include DeepSeek-Coder-MQA-5.7B Guo et al. (2024), DeepSeek-Coder-base-6.7B Guo et al. (2024), Qwen2.5-Coder-3B-base Hui et al. (2024), Qwen2.5-Coder-7B-base Hui et al. (2024) and Qwen2.5-Coder-14B-base Hui et al. (2024).
Common model. We evaluate commerical models including GPT-5 Leon (2025), Gemini 2.5 Pro Comanici et al. (2025), Grok4 Adinath and IS (2025), and Deepseek V3.2 Liu et al. (2025a).
Metrics. We compare the generated code with the reference and compute the exact match (EM), edit similarity (ES) and Bilingual Evaluation Understudy (BLEU) following Liu et al. (2025b); Zhang et al. (2023). Their detailed implementations are defined in A.3
Model Chisel V/SV VHDL HLS
EM ES BLEU EM ES BLEU EM ES BLEU EM ES BLEU
Qwen2.5-Coder-3B 0.0 7.9 5.4 0.2 7.9 2.8 0.0 6.6 2.9 0.0 5.5 4.1
+Tuning 24.8 54.3 50.6 34.8 65.5 56.5 28.2 59.1 60.2 33.1 55.0 55.7
Qwen2.5-Coder-7B 0.0 14.8 4.9 0.0 12.7 3.3 0.0 13.0 3.8 0.2 14.1 4.5
+Tuning 26.6 66.7 56.6 39.0 70.8 60.1 34.5 71.9 64.9 35.5 69.5 58.3
Qwen2.5-Coder-14B 0.0 16.5 6.9 1.0 18.5 7.5 0.4 15.8 5.7 1.1 16.0 6.3
+Tuning 31.3 68.9 59.6 41.9 74.4 64.2 35.3 73.8 67.7 38.2 71.5 61.3
DS-Coder-5.7B 0.0 13.2 3.7 0.4 8.8 2.6 0.0 9.0 3.2 0.0 9.7 2.8
+Tuning 24.8 61.2 50.1 37.6 71.7 61.0 30.7 71.1 63.0 33.1 67.7 55.9
DS-Coder-6.7B 0.0 5.4 3.4 0.2 5.7 2.5 0.0 5.6 2.6 0.0 3.6 2.2
+Tuning 25.7 56.5 52.6 37.8 67.4 61.2 31.5 61.1 63.9 36.8 60.4 58.2
Table 4: Exact Match (EM %), Edit Similarity (ES %), and BLEU scores % across different hardware description languages after fine-tuning for code LLMs.
Model Retrieval Method Chisel V/SV VHDL HLS
EM ES BLEU EM ES BLEU EM ES BLEU EM ES BLEU
GPT-5 No RAG 24.8 56.7 45.9 34.5 60.9 49.5 15.4 52.0 41.7 31.4 63.7 55.3
RepoCoder 27.6 59.6 49.9 35.4 61.3 50.2 16.7 54.1 43.5 34.4 63.9 55.5
GraphCoder 13.9 44.2 36.4 15.7 40.1 31.8 7.5 35.7 29.0 24.9 49.5 41.1
RLCoder 27.8 60.2 50.3 36.2 61.7 51.4 17.4 55.3 45.7 36.9 72.9 60.1
Gemini 2.5 Pro No RAG 33.2 59.9 49.8 33.2 59.9 49.8 16.3 50.8 41.4 34.6 63.8 57.3
RepoCoder 33.7 61.7 54.7 34.8 59.6 50.5 18.0 52.3 42.3 39.4 65.3 59.1
GraphCoder 13.3 50.8 44.0 22.2 50.5 43.2 9.1 45.0 39.8 29.1 57.8 50.8
RLCoder 34.6 63.2 55.1 35.3 60.2 51.3 19.2 53.7 43.5 41.2 66.9 60.2
Grok4 No RAG 22.4 53.7 42.1 31.4 56.8 46.7 14.3 47.8 36.3 31.6 60.5 52.3
RepoCoder 23.8 54.7 44.3 32.0 56.1 45.5 15.7 49.3 37.4 32.3 61.1 53.7
GraphCoder 12.7 35.4 30.5 16.5 38.1 31.8 8.1 35.3 30.3 18.0 38.8 33.1
RLCoder 25.6 55.2 43.1 33.4 57.2 47.9 16.8 51.2 38.2 33.3 62.2 61.3
Deepseek V3.2 No RAG 22.6 51.2 39.8 28.7 54.1 42.2 13.3 43.0 30.0 30.3 57.3 49.5
RepoCoder 23.2 52.1 41.6 28.3 52.6 40.7 14.2 42.8 29.2 33.3 57.9 50.8
GraphCoder 9.3 39.5 32.8 15.1 39.1 32.1 9.9 39.1 33.0 22.7 47.9 40.3
RLCoder 24.3 53.1 40.2 29.4 45.1 42.4 14.2 43.6 31.3 34.3 59.2 51.0
Table 5: EM %, ES %, and BLEU scores % with different retrieval methods across hardware description languages.
4.2 Experimental Setup
(a) Impact of dataset size on EM performance.
(b) Impact of dataset size on ES performance.
(c) Model performance and the number of lines to be completed.
Figure 4: Impact of training data scale and completion length on hardware code completion.
(a) EM of GPT-5 without retrieval augmentation (No-RAG) under increasing prompt lengths.
(b) EM of post-trained Qwen2.5-Coder-7B under increasing prompt lengths.
(c) Model performance and code structure depth.
Figure 5: Effects of context length and code structure depth on hardware code completion.
(a) EM performance of different categories.
(b) ES performance of different categories.
Figure 6: Performance of GPT-5 and post-trained Qwen2.5-Coder-7B across hardware semantic categories.
Baseline. We provide the model only with the file that contains the completion cursor. Without explicit inter-file context, the model must depend on its pretrained ability to produce the completion.
Tuning. To further improve the performance of repository-level code completion, we fine-tune code LLMs on training dataset supervised fine-tuning (SFT) with detailed configuration in Appendix A.3.
Retrieval Framework. There are many different repository-level code completion methods including RepoCoder (Zhang et al., 2023), GraphCoder (Liu et al., 2024b), and RLCoder Wang et al. (2024). We evaluate these frameworks with different models on MHRC-Bench-Eval. The total input window is fixed at 2,048 tokens for retrieved context and local context for all the models.
4.3 Main Results
Table 4 shows the evaluation results of base models and fine-tuned models in different hardware languages. Prior to supervised fine-tuning (SFT), all models achieve nearly zero EM scores, reflecting the difficulty of repository-level hardware code completion for base models. After fine-tuning on our dataset, the performance of all models improves substantially, with fine-tuned variants significantly outperforming their base models. Overall, the post-trained Qwen2.5-Coder-14B consistently outperforms Qwen2.5-Coder-7B, which in turn achieves better performance than DS-Coder-6.7B and DS-Coder-5.7B, and substantially surpasses Qwen2.5-Coder-3B. We detail how SFT improves model performance in Appendix A.4.
Table 5 reports the evaluation results of different retrieval methods. With the exception of GraphCoder, all retrieval approaches consistently improve model performance across multiple hardware languages. The reasons for GraphCoder’s inferior performance are discussed in Appendix A.5. Among the evaluated retrieval methods, RLcoder achieves the strongest overall performance. When comparing models using the same retrieval strategy, Gemini 2.5 Pro attains the best overall results, while GPT-5 exhibits better performance on specific metrics, including EM and BLEU for V/SV, as well as ES and BLEU for VHDL.
Table 6: Models’ CodeBLEU (%) across hardware description languages. CodeBLEU calculation is introduced in Appendix A.6.
Model Chisel V/SV VHDL HLS
Qwen2.5-Coder-7B 14.8 22.9 22.2 21.0
+ Tuning 49.3 54.2 56.3 52.7
GPT-5 (No RAG) 42.7 45.9 41.2 51.0
Comparison between pretrained and fine-tuned models show that fine-tuned models consistently achieve higher EM, ES, and BLEU scores than larger, general-purpose models with substantially more parameters. For example, Qwen2.5-Coder-14B, after fine-tuning on our dataset, outperforms GPT-5, Grok-4, and DeepSeek V3.2 across all hardware languages, and surpasses Gemini 2.5 Pro in all cases except Chisel. These results demonstrate that our training dataset effectively enhances the hardware code completion capability of code LLMs, even with a smaller model scale.
4.4 Experimental Analysis
4.4.1 Analysis on CodeBLEU and Compilation rate.
CodeBLEU (Ren et al., 2020) evaluates code generation quality by considering not only surface-level n-gram overlap but also syntactic parsability and semantic consistency. In Table 6, we report CodeBLEU scores across four hardware description languages. In addition, Table 7 presents the compilation success rates for different languages. Together, these results show that fine-tuning substantially improves the performance of Qwen2.5-Coder-7B.
Table 7: Models’ compilation rate (%) across hardware description languages. The procedure for computing the compilation rate is described in Appendix A.7.
Model Chisel V/SV VHDL HLS
Qwen2.5-Coder-7B 0.0 0.2 0.0 0.2
+ Tuning 28.8 40.7 35.2 37.7
GPT-5 (No RAG) 27.0 36.4 18.9 34.6
4.4.2 Analysis on different training data sizes.
In Figure 4, we study the effect of training dataset size on model performance. Specifically, we randomly sample 25%, 50%, 75%, and 100% of the training data to fine-tune Qwen2.5-Coder-7B-base. Evaluation on the test set shows a clear and consistent performance improvement as the training data size increases for most languages, indicating that larger and more informative datasets lead to stronger completion capabilities. An exception is observed for HLS from 75% to 100%, which is closely related to C/C++. The pretrained model already exhibits strong prior knowledge in HLS-style C/C++ code. Additional fine-tuning data may introduce distributional shifts that interfere with this prior, resulting in negative transfer. Overall, these results further validate the effectiveness and scalability of our dataset.
4.4.3 Analysis on different lines.
As shown in Figure 4(c), post-trained Qwen2.5-Coder-7B and GPT-5 can effectively complete tasks involving a small number of lines. However, as the number of lines to be completed increases, the scores of the generated code gradually decline. This indicates that completing multi-line code remains a challenge for code LLMs.
4.4.4 Analysis on the prompt length.
For a fair comparison across models, we set the default prompt length to 2048 tokens. Nevertheless, varying the prompt length leads to different performance trends for both fine-tuned and general-purpose models, as illustrated in Figure 5. We therefore evaluate prompt lengths of 512, 1024, 2048, and 4096 tokens. The results show that GPT-5 achieves consistently higher EM performance at 2048 and 4096 tokens than at shorter prompt lengths. In contrast, the fine-tuned Qwen2.5-Coder-7B performs better at 2048 tokens than at 4096 tokens. Based on these observations, we adopt 2048 tokens as the input prompt length, which provides strong and stable performance across models.
4.4.5 Analysis on code structural levels.
As mentioned in 3.5, we categorize our test cases into five types according to node depths in the concrete syntax tree. As shown in Figure 5(c), We present the performance of post-trained Qwen2.5-Coder-7B-base and GPT-5 on the test set across these different structural levels, and we observe that as the structural level decreases, the performance of LLM may either improve or degrade depending on the structure granularity, which is different from software shown in Liu et al. (2025b) where performance decreases consistently with increasing structural levels. For detailed performance across node types, please refer to Appendix A.8.
4.4.6 Analysis on hardware semantic levels.
Mentioned in section 3.5, we categorize the nodes within the concrete syntax tree into nine primary semantic levels based on their semantic characteristics, and we provide the performance of post-trained Qwen2.5-Coder-7B-base and GPT-5 for these various semantic levels across multilingual languages on the test set, as illustrated in Figure 6(a) and 6(b). Notably, we observe significant performance disparities across different semantic levels. Specifically, models show superior performance on “Declaration and Definition”, while exhibiting lower efficacy on “Control Flow Block”. This suggests that current code LLMs are proficient at completing tasks related to variable definitions and references, yet their capacity to handle complex codes are limited. Appendix A.9 details performance on different language’s semantic types.
4.5 Key Observations
Performance varies significantly across hardware languages. Model performance differs substantially across hardware description languages. SystemVerilog/Verilog and HLS consistently achieve the best performance across all evaluated LLMs, whereas VHDL exhibits the lowest. This gap highlights the uneven maturity of LLMs in handling different hardware design paradigms.
Post-training is critical for improving hardware code completion. Post-training substantially enhances LLM performance on hardware code completion tasks. After fine-tuning on MHRC-Bench-train, code models show significant improvements over their base models and are able to surpass several general-purpose models. This demonstrates the effectiveness of targeted post-training for hardware-oriented tasks.
Similarities between hardware and software code: (1) Increasing the target length consistently degrades performance for both hardware and software code completion tasks. (2) LLMs perform well on identifier-level generation in both domains, indicating strong lexical and naming capabilities.
Differences between hardware and software code. (1) Structural sensitivity: For software code, deeper concrete syntax tree (CST) structures often correlate with improved LLM performance Liu et al. (2025b). In contrast, for hardware code, increasing structural complexity typically leads to performance degradation. (2) Prompt length sensitivity. Increasing prompt length generally benefits software code completion, whereas for hardware code it yields limited or inconsistent gains (from 2048 to 4096 tokens), reflecting the greater difficulty of reasoning about hardware semantics.
Inappropriate retrieval can degrade performance. Retrieval-augmented methods are not always beneficial. We observe that retrieving irrelevant or noisy information can harm model performance, as exemplified by GraphCoder. This highlights the importance of retrieval quality and task-aware retrieval strategies.
Conclusion
We present MHRC-Bench, the first multilingual, repository-level benchmark for hardware code completion, covering RTL, HLS, and generator-based hardware languages across both design and verification stages. By introducing fine-grained annotations at the code-structure and hardware-semantic levels, MHRC-Bench enables detailed analysis of LLM behavior in hardware-oriented settings. Extensive experiments show that post-training on MHRC-Bench-Train substantially improves hardware code completion performance, allowing relatively small open-weight models to outperform strong commercial baselines. Our findings highlight fundamental differences between hardware and software code completion and establish MHRC-Bench as a valuable resource for advancing LLMs in hardware design automation.
Limitations
First, while MHRC-Bench covers multiple major hardware description paradigms, it does not yet include analog/mixed-signal or low-level physical design languages. Second, our evaluation focuses on textual similarity metrics, which may not fully capture functional correctness of generated hardware code. Third, MHRC-Bench does not cover analog/mixed-signal and physical design languages but mainly digital hardware description languages. Addressing these limitations is an important direction for future work.
Ethics Statement
MHRC-Bench is constructed entirely from publicly available GitHub repositories released under permissive open-source licenses (e.g., MIT, BSD, Apache). Repositories with restrictive or copyleft licenses are excluded, and duplicate repositories and vendor/dependency directories are removed. The dataset contains only source code and does not intentionally include personal or private information. MHRC-Bench covers multiple hardware languages and abstraction levels to reduce evaluation bias, though it reflects the distribution of open-source hardware projects and may not represent proprietary practices. The benchmark is intended solely for research on hardware code completion and model evaluation. Generated code should not be deployed without proper human review, simulation, and verification.
While the proposed system automates aspects of hardware design, we recognize that AI-generated design suggestions should always be validated by domain experts before deployment in safety-critical or commercial environments. We encourage responsible use and open scientific collaboration. Also, we use ChatGPT to polish our paper writing.
Appendix A Appendix
A.1 Cross-file dependency
The target file often depends on functions, definitions, or modules defined in other files to be completed correctly. Since different hardware description languages follow distinct grammatical and structural rules, we adopt language-specific strategies to identify and extract cross-file dependencies.
For each task file, we compute cross-file dependencies as the number of distinct other files within the same repository that the file depends on, using deterministic, language-specific resolution rules. For HLS C/C++, dependencies are inferred from resolved include directives, macro references, and user-defined type usages. For VHDL, dependencies are determined by references to design units such as entities, packages, and contexts. For Chisel/Scala, dependencies are identified through resolved imports and statically resolvable symbol usages. For SystemVerilog/Verilog, dependencies are computed from includes, package references, module or interface instantiations, macro usage, and typedef references. In all cases, the final dependency count equals the number of uniquely resolved dependent files, excluding the file itself, and results are reported as averages across task items (or aggregated per file or repository, depending on the evaluation setting).
A.2 Distribution across hardware sematic categories
Based on the hardware semantic categories defined in Table 3, we categorize completion targets in each hardware description language accordingly and present their distributions in Figure 7.
(a) Chisel
(b) HLS
(c) SystemVerilog
(d) VHDL
Figure 7: Category distribution of completion targets in MHRC-Bench-eval. The distributions reveal clear language-specific patterns, reflecting differences in hardware design abstractions, coding styles, and verification practices across Chisel, HLS, SystemVerilog/Verilog, and VHDL.
A.3 Evaluation Configurations
A.3.1 Prompts
The exact prompt used during model inference is given in Figure 8.
Task Prompt You are a precise code-completion model completing a repository blank completion task. Input: a repository concatenation where each file of the repo is wrapped in --- FILE: <repo_name>/<rel/path/to/file> (BEGIN) ---<file content...> --- FILE: <repo name>/<rel/path/to/file> (END) ---. Some lines have been replaced by the exact token <TARGET>. Your task: reconstruct only the replaced line, using the repository context. Note that for certain large repositories, functions of lesser importance were removed of their implementation, and only the signature/documentation were kept for the sake of simplicity.
OUTPUT: the reconstructed line(and nothing else).
<Left Context>... ...</Left Context>
<TARGET>
<Right Context>... ...</Right Context>
Figure 8: The prompt used for repository-level code completion (No RAG). If the context length exceed the token limit, the head of the left context and the tail of the right context are trimmed.
A.3.2 Fine-Tuning
For fine-tuning local LLMs (Qwen and Deepseek), the training objective was modeled as a "fill-in-the-middle" instruction task, where the model is prompted with a concatenated repository context containing file delimiters and a specific <TARGET> token representing the missing line. To manage context window constraints, inputs were dynamically truncated to a maximum of 2,048 tokens, employing a balanced budgeting strategy to prioritize the immediate left and right context surrounding the target. Fine-tuning was achieved using Low-Rank Adaptation (LoRA) applied to all linear attention and feed-forward layers, configured with a rank $r=16$, scaling factor $\alpha=32$, and a dropout of $0.05$. The model was optimized using AdamW with a peak learning rate of $2\times 10^{-4}$, following a cosine decay schedule with a 3% warm-up period all over a single epoch. Cross-entropy loss was computed exclusively on the generated ground-truth line, masking the instruction and context tokens. For reference, we plot the loss against the epoch number for Qwen2.5-Coder-7B, Qwen2.5-Coder-14B, and deepseek-coder-5.7bmqa-base in Figure 9.
A.3.3 Evaluation Metrics
During inference, the model was configured with a temperature of $0.0$, a top_p of 1, and a maximum generation length of $64$ tokens. We choose a maximum generation length of 64 tokens because all ground-truth completions in our dataset contain no more than 64 tokens. Generation is stopped when the eos_token was produced or when the maximum length was reached.
Prior to calculating the EM, ES and BLEU metrics, the generated code were stripped of comments, followed by trimming whitespace and dropping empty lines. The EM is set to $1$ if the cleaned generation results exactly matches the ground truth, otherwise it is set to 0. For ES, we evaluate the following formula:
$\mathrm{ES}\;=\;1-\frac{\operatorname{editdistance}\!\left(\text{target\_str},\text{prediction\_str}\right)}{\max\!\left(1,\;\lvert\text{target\_str}\rvert,\;\lvert\text{prediction\_str}\rvert\right)}$
where editdistanceeditdistance is the Levenshtein edit distance between the generated code and the ground truth. $BLEU$ is calculated as the token-level BLEU between the cleaned generation results and the ground truth.
A.3.4 Commercial API setting
For evaluating commercial LLMs, we utilized API services provided by ChatfireAPI. The specific model names are gpt-5, deepseek-v3.2, gemini-2.5-pro, and grok-4. API defaults were used for temperature, top_p, max_tokens, etc.
Figure 9: Training loss over epochs for Qwen2.5 and DeepSeek models.
A.4 Case Study of Pretrained-Model Responses
To illustrate the effect of fine-tuning on local models, we analyze the completions generated by Qwen2.5-Coder-7B before and after fine-tuning. Specifically, we list for each tested language common errors made by the model that were significantly improved by fine-tuning.
A.4.1 Verilog/SystemVerilog
Syntax Hallucination: One of the most frequent mistakes made by the pre-trained model is completely losing track of the syntax context. in AXI-SDCard-High-Speed-Controller, for instance, the model hallucinates a module instantiation statement .waddr(waddr),.wdata(wdata)... instead of the expected IO declaration input logic[2:0] axilite_arprot
Structural Hallucination: The pre-trained model is prone to regurgitating file headers, copyright notices, or artificial file boundary markers in the middle of code blocks. In custom_uvm_report_server for instance, the pre-trained model suddenly outputs a full copyright header // Copyright 2007-2010 Mentor Graphics...inside the build_phase function instead of the expected super.build_phase(phase); call.
Pattern Oversight: The pre-trained model fails to follow obvious naming sequences or structural patterns established immediately before the target line. In SystemVerilog Assertions, for instance, the model failed to recognize the naming pattern of logic gnt1; logic gnt2; and even missed the comment //3 Requests to Arbiter logic req1;logic req2; logic req3;. Instead, it regurgitated earlier comments and produced irrelevant code.
Logical Confusion: The pre-trained model sometimes generates code that is syntactically valid SystemVerilog but logically disconnected from the immediate control flow (e.g., jumping to a different state in a Finite State Machine). In virtio, the code is strictly defining an always_ff block for reset logic. The pre-trained model attempts to insert FSM state logic _EVENT_IDX: begin... from a case statement that belongs elsewhere, ignoring the always_ff structure.
A.4.2 High-Level Synthesis
Structural Hallucination: Similar to other languages, the pre-trained model for HLS code often completely loses track of the file position and hallucinates the start of a new file. Instead of generating the next line of code, the model outputs copyright notices, include guards, or library imports. For example, in hls-spmv, the pre-trained model outputs #include "spmv.h" despite the cursor being positioned in the spmv_kernel function.
C++/HLS Confusion: HLS often requires specific syntax not included in standard C++ that are used for reporting and optimization directives. The pre-trained model often writes valid standard C++ but misses these hardware-description necessities. In hls_spmv, for instance, the ground truth required an L1: in the loop L1: for (int i = 0; i < NUM_ROWS; i++), which the pre-trained model neglected.
Syntax Hallucination: Similar to other languages, the pre-trained model fails to recognize the contextual logic and syntax. In libllsm, for instance, while the context explicitly requires a memory allocation (calloc), the pre-trained model instead attempts to define a new function FP_TYPE* llsm_envelope....
A.4.3 VHDL
Structural Confusion: Similar to other languages, the pre-trained model struggles to comprehend the context and has the tendency to hallucinate file headers, library declarations, or copyright notices in the middle of an active code block. For the FPGA-Speech-Recognition repository, for instance, the pre-trained model outputted library IEEE; use IEEE.STD_LOGIC_1164.ALL;... entity UART_PARITY.... despite the target line being deep inside a generate block for odd parity. More severely, the model often confuses code segments with prompt structures and output strings used to mark file boundaries in prompts. In the cpu_for_nscscc2022_single repository, the model outputs --- FILE: dcache_tag.vhd (END) --- immediately, skipping the required generic map assignments C_MEM_TYPE => 1.
Syntax Confusion: VHDL distinguishes heavily between concurrent statements and sequential statements inside processes or subprograms. The pre-trained model often mixes these contexts. In FPGA-I2C-Minion, the pre-trained model hallucinates unrelated procedure calls i2c_wait_quarter_clock or definitions that belong in a testbench or architecture body, not the sequential spike generator logic that the target line is in.
A.4.4 Chisel
Structural Confusion: Similar to other languages, the pre-trained model often hallucinates copyright headers and library declarations.
Language Confusion: The pre-trained model sometimes fails to generate valid Chisel syntax or uses generic tokens that don’t fit the Chisel hardware construction paradigm. For example, in rjrouter, the ground truth is c => new GCDUnitTester(c), but the pre-trained model outputs <|fim_middle|>, indicating a complete failure to generate content.
A list of examples referenced above can be found in Table 9
(a) HLS: Line Number
(b) HLS: Structural Complexity
(c) SystemVerilog: Line Number
(d) SystemVerilog: Structural Complexity
Figure 10: EM performance for HLS and SystemVerilog. Top-to-bottom: line number then structural complexity.
(a) Chisel: Line Number
(b) Chisel: Structural Complexity
(c) VHDL: Line Number
(d) VHDL: Structural Complexity
Figure 11: EM performance for Chisel and VHDL. Top-to-bottom: line number then structural complexity.
A.5 GraphCoder Modifications
The original implementation of GraphCoder is designed for left-to-right completions using only the left context. To fit our objective of fill-in-the-middle completions, we modified GraphCoder such that the right context is also preserved and included during prompt assembly. As the graph builder and the retriever operate independently of the left and right contexts, this has no other effect on the functionality of GraphCoder.
Furthermore, the original implementation of GraphCoder included helper functions in utils/ccg.py for traversing the CST that only supported Java and Python. To extend support to hardware languages, we added helper functions to handle each language’s unique syntax and identify control-flow and data-dependence. Other minor changes were made in various locations to include HDL languages during tree-sitter Parser language selection and code file filtering.
We note that GraphCoder under-performs on HDL repositories when compared with other methods. We attribute this to the Retriever, which demonstrates lower hit@k metrics in HDL languages than in software languages like Python. This is illustrated in Table 8. By GraphCoder’s design, ineffective retrieval will flood the context window with irrelevant code snippets, thus causing GraphCoder to perform worse than No RAG.
Language Hit@1 Hit@5
Scala 0.03 0.05
VHDL 0.11 0.16
SystemVerilog 0.09 0.12
HLS 0.19 0.25
Python (API-level) 0.29 0.35
Python (Line-level) 0.27 0.32
Table 8: Retriever hit@k on HDL languages and Python
A.6 CodeBLEU
CodeBLEU Metric.
We adopt a practical, language-aware variant of CodeBLEU to evaluate hardware code generation across Chisel, SystemVerilog, VHDL, and HLS (C/C++). The final score is computed as a weighted combination of four complementary components:
$\text{CodeBLEU}=$ $\alpha\cdot\text{BLEU}+\beta\cdot\text{WeightedBLEU}$ (1)
$+\gamma\cdot\text{SyntaxMatch}$
$+\delta\cdot\text{DataflowMatch}.$
where we set $\alpha=\beta=\gamma=\delta=0.25$ in all experiments.
Token-level BLEU.
We compute standard token-level BLEU after language-aware normalization, including comment removal and whitespace normalization. Code is tokenized into identifiers, keywords, literals, and operators. Smoothing is applied to ensure stability on short code snippets.
Weighted BLEU.
To emphasize language-specific keywords and hardware constructs, we employ a weighted BLEU variant. Tokens corresponding to critical constructs (e.g., always_ff, process, when, RegInit, ap_uint) are assigned higher weights by proportionally expanding their occurrences during BLEU computation. This component encourages the preservation of semantically important syntax.
Syntax Match.
We measure syntactic similarity using Tree-sitter–based parsing. Each snippet is minimally wrapped to ensure parsability, and we extract a multiset of concrete syntax tree (CST) node types. The syntax score is computed as the multiset F1 score between the predicted and reference node-type distributions, capturing structural similarity beyond surface tokens.
Dataflow Match.
To assess semantic consistency, we approximate dataflow using heuristic def–use analysis. Language-specific assignment patterns are used to extract variable dependency edges, and the dataflow score is computed as the F1 score between the predicted and reference dependency sets. This component rewards correct preservation of variable usage relationships even when exact syntax differs.
Discussion.
Together, these components balance lexical accuracy, keyword fidelity, structural correctness, and semantic consistency, making the metric well suited for repository-level hardware code completion, where exact string matching alone is insufficient.
A.7 Compilation rate calculation.
Compilation is evaluated per completed snippet by embedding it into a minimal, deterministic wrapper to ensure syntactic completeness, followed by a single-pass syntax or compilation check with a fixed timeout. A snippet is considered successfully compiled if the tool terminates without syntax or elaboration errors. The compilation rate is computed as the number of successfully compiled test cases divided by the total number of test cases for each language.
For SystemVerilog/Verilog, we use Verilator (v5.x) with --lint-only, wrapping snippets in a minimal module context when necessary. For VHDL, we use GHDL (v3.x) with the VHDL-2008 standard, wrapping snippets in a minimal entity and architecture. For Chisel, we use the Scala compiler (Scala 2.12) with Chisel3, wrapping snippets in a minimal Scala object containing a Module definition and performing frontend compilation only. For HLS C/C++, we use clang (v14+) in syntax-checking mode (-fsyntax-only), wrapping snippets in a dummy top-level function when necessary.
A.8 Detailed Performance on line numbers and structural levels
We present how the performance of Qwen2.5-Coder-7B and RepoCoder (Driven by GPT-5) varies with the number of lines to be completed and structural complexity for different hardware languages in Figures 10 and 11.
A.9 Detailed Performance on hardware semantic levels.
We present how the performance of Qwen2.5-Coder-7B and RepoCoder (Driven by GPT5) varies with different hardware semantic levels for different hardware language.
(a) Chisel: EM by Category
(b) Chisel: ES by Category
(c) Chisel: BLEU by Category
Figure 12: Per-category performance on Chisel. Metrics are reported in terms of EM, ES, and BLEU.
(a) SystemVerilog: EM by Category
(b) SystemVerilog: ES by Category
(c) SystemVerilog: BLEU by Category
Figure 13: Per-category performance on SystemVerilog across EM, ES, and BLEU.
(a) HLS: EM by Category
(b) HLS: ES by Category
(c) HLS: BLEU by Category
Figure 14: Per-category performance on HLS code completion.
(a) VHDL: EM by Category
(b) VHDL: ES by Category
(c) VHDL: BLEU by Category
Figure 15: Per-category performance on VHDL across different evaluation metrics.
Figure 16: Example one of chisel code completion examples
Figure 17: Example two of chisel code completion examples
Figure 18: Example one of HLS code completion examples
Figure 19: Example two of HLS code completion examples
Figure 20: Example two of systemverilog code completion examples
Figure 21: Example two of systemverilog code completion examples
Figure 22: Example two of VHDL code completion examples
Figure 23: Example two of VHDL code completion examples
A.10 Code Completion Examples
For each hardware description language, we select two representative repositories as examples to illustrate cross-file context, in-file context, the selected completion target, and the corresponding ground truth in the following figures.
Lang. Repo Name Ground Truth Pre-Train Output (Failure)
Verilog AXI-SDCard-High-Speed-Controller input logic[2:0] axilite_arprot, .waddr(waddr), .wdata(wdata), .we(we)... (Hallucinated Instantiation)
Verilog custom_uvm_report_server super.build_phase(phase); // Copyright 2007-2010 Mentor Graphics... (Header Regurgitation)
Verilog SystemVerilogAssertions logic gnt3; // Write assertions for a simple Round Robin... (Context Reset)
Verilog virtio always_ff @(posedge aclk...) _EVENT_IDX: begin tx.tvalid <= 1’b1; ... (Logic Hallucination)
HLS hls-spmv int col_left = 0; #include "spmv.h" #include <hls_stream.h>... (Header Regurgitation)
HLS hls-spmv L1: for (int i = 0; ... void spmv(int rowPtr[NUM_ROWS+1]... (Function Definition)
HLS libllsm FP_TYPE* x = calloc(ns... FP_TYPE* llsm_envelope(FP_TYPE* spectrum... (Function Definition)
VHDL FPGA-Speech-Recognition PARITY_OUT <= parity_temp; library IEEE; use IEEE.STD_LOGIC_1164.ALL;... (Header Regurgitation)
VHDL cpu_for_nscscc2022_single C_HAS_WE => 1, C_MEM_... --- FILE: dcache_tag.vhd (END) --- (Premature EOF)
VHDL FPGA-I2C-Minion if ENDSIM = false then... i2c_wait_quarter_clock; scl_pre_spike <= ’Z’; (Unrelated Logic)
Chisel rjrouter class Lfsr16Test extends... <|fim_middle|> (Model Failure Token)
Table 9: Comparison of Ground Truth vs. Pre-Training Failures across Hardware Languages.

### v2

@@
 We conduct a comprehensive evaluation of models on MHRC-Bench-Eval.
-Comprehensive evaluation results and analysis demonstrate the effectiveness of MHRC-Bench.
-MHRC-Bench: A Multilingual Hardware Repository-Level Code Completion benchmark
-Qingyun Zou1, Jiahao Cui1, Nuo Chen1, Bingsheng He1, Weng-Fai Wong1 1National University of Singapore qingyunzou@u.nus.edu
-Figure 1: Average EM and ES comparison of different LLMs on the MHRC-Bench-Eval benchmark.
+Our evaluation reveals insights into the differing behaviors of LLMs in hardware code completion across multiple hardware languages.
+Machine Learning, ICML
 Figure 1: Overview of our hardware code completion benchmark across multiple hardware description paradigms.
-We present three representative examples from generator-based hardware construction languages (e.g., Chisel), high-level synthesis (HLS) languages (e.g., Xilinx HLS), and Register-Transfer Level (RTL) design languages (SystemVerilog).
+We present three representative examples from domain-specific language (DSL) (e.g., Chisel), high-level synthesis (HLS) languages (e.g., Xilinx HLS), and Register-Transfer Level (RTL) design languages (SystemVerilog).
 For each example, both in-file and cross-file contexts are provided, with the completion target explicitly marked.
@@
 1 Introduction
-Recent advancements in large language models (LLMs) have led to significant progress in code-related tasks such as code understanding, generation, and completion (Liu et al., 2023; Nam et al., 2024; Liu et al., 2024a; Pan et al., 2024; Raihan et al., 2025; Yang et al., 2024).
-These models, trained on vast datasets encompassing both natural language and source code, have demonstrated remarkable proficiency in modeling programming languages as structured sequences and predicting logical code continuations (Liang et al., 2024; Li et al., 2025; Guo et al., 2024; Hui et al., 2024; Liu et al., 2024b; Abedu et al., 2025; Hua et al., 2025).
+Recent advancements in large language models (LLMs) have led to significant progress in code-related tasks such as code understanding, generation, and completion (Liu et al., 2023b; Nam et al., 2024; Liu et al., 2024a; Pan et al., 2024; Raihan et al., 2025; Yang et al., 2024).
+These models, trained on vast datasets encompassing both natural language and source code, have demonstrated remarkable proficiency in modeling programming languages as structured sequences and predicting logical code continuations (Liang et al., 2024; Li et al., 2025; Guo et al., 2024; Hui et al., 2024; Liu et al., 2024c; Abedu et al., 2025; Hua et al., 2025).
 Consequently, LLM-based code completion has emerged as a valuable tool in real-world software development environments.
@@
 (2025) equip LLMs with external tool APIs to more effectively utilize repository-level context during the code completion process.
-While recent work has explored Verilog code completion (Wu et al., 2025b; Allam and Shalan, 2024), existing benchmarks remain narrowly focused on a single hardware language.
-Inspired by Liu et al.
-(2025b), we propose MHRC-Bench, the first multilingual repository-level benchmark for hardware code completion, consisting of MHRC-Bench-Train and MHRC-Bench-Eval.
+Despite their success in software code, LLMs’ performance on hardware code remains unclear, as it differs fundamentally in execution semantics, structural parallelism, and cross-file dependency patterns.
+In particular, hardware code completion requires understanding static structure, signal connectivity, and design intent rather than purely sequential execution.
+We introduce MHRC-Bench including MHRC-Bench-Train and MHRC-Bench-Eval, a multilingual repository-level benchmark designed to expose structural and semantic failure modes of LLMs in hardware code completion.
 Using the MHRC-Bench-Train dataset to fine-tune code LLMs, we observe that billion-scale fine-tuned models such as Qwen2.5-Coder-7B achieve substantially higher EM and ES scores than their pretrained counterparts, whose EM is nearly zero on hardware code completion tasks.
-After fine-tuning, these models also outperform larger general-purpose models, including GPT-5, Grok-4, Gemini 2.5 Pro, and DeepSeek V3.2, as shown in Figure 1.
+After fine-tuning, the post-trained Qwen2.5-Coder-7B also outperform larger general-purpose models such as GPT-5 about EM, ES, and compilation pass rate.
+with the same input window.
 This demonstrates that MHRC-Bench-Train effectively enhances the hardware code completion capabilities of smaller-scale code LLMs.
@@
 At the hardware-semantic level, code is categorized into nine subcategories, enabling detailed evaluation across diverse hardware semantics.
-The benchmark and related code are publicly available. The contributions are summarized as follows:
+The benchmark, related code and fine-tuned models are publicly available. The contributions are summarized as follows:
   • We introduce MHRC-Bench-Eval, the first large-scale multilingual repository-level hardware code completion benchmark, covering three major hardware programming paradigms across the hardware design flow.
@@
   • We construct MHRC-Bench-Train, a training dataset designed to enhance repository-level hardware code completion.
-Fine-tuning on MHRC-Bench-Train leads to substantial performance gains, even enabling smaller open-weight models to outperform strong commercial baselines.
-  • We conduct comprehensive evaluations across a wide range of models, including commercial LLMs and open-source code LLMs, as well as state-of-the-art retrieval-based methods.
+Fine-tuning on MHRC-Bench-Train leads to substantial performance gains, even enabling smaller open-weight models to outperform larger general models.
+  • We conduct comprehensive evaluations across a wide range of models, including general LLMs and open-source code LLMs, as well as state-of-the-art retrieval-based methods.
 Our analysis reveals key performance trends and insights into LLM behavior in hardware code completion.
-2 Related Works and Background
+Table 1: Comparison with hardware code benchmark.
+FG indicates fine-grained annotation.
+Train means if it contains training dataset.
+NL means natural language.
+RTL means register transfer level.
+HLS means High-Level Synthesis.
+DSL means domain-specific language.
+Benchmark Level Task Type Input Languages FG Train Output Data types Samples
+VerilogEval Module Code Gen NL Spec RTL × Module Code Synthetic and Real-world 8,502
+RTLLM Module Code Gen NL Spec RTL × × Module Code Real-world 30
+HLS-Eval Module Code Gen NL Spec or C Code HLS × × Module Code Real-world 94
+HDLEval Module Code Gen NL Spec RTL, DSL × × Module Code Real-world 461
+RTL-Repo Repository Code Completion Repository Code RTL × Single-line Real-world 4098
+Ours Repository Code Completion Repository Code RTL,HLS,DSL Single and Multi-line Real-world 47,175
+2 Related Works
 2.1 Repository-level Code Completion.
-Recent repository-level code completion methods (Deng et al., 2025; Li et al., 2025; Liu et al., 2024b; Liang et al., 2024; Liu et al., 2025d; Wu et al., 2025b; Wang et al., 2024; Zhang et al., 2025) aim to retrieve related code snippets across files within a repository, where existing datasets mainly focus on software engineering task.
-For example, M2RC-Eval (Liu et al., 2025b) supports Java, Python and so on these related software codes completion.
-2.2 LLM for EDA.
-Recent advancements in applying large language models (LLMs) to electronic design automation (EDA) tasks, such as RTL/HLS code generation and hardware verification, have shown significant promise.
-Notable studies in NLP venues include SynC-LLM (Liu et al., 2025c), which uses diffusion models and LLM prompting to generate realistic digital circuits, improving downstream EDA model performance.
-Wu et al.
-(2025a) introduces a multi-agent LLM system that enhances the reliability of multi-step EDA tool flows.
-DeepRTL2 (Liu et al., 2025e) showcases LLMs’ versatility in RTL generation and embedding-based tasks, setting new benchmarks.
-Lastly, AssertionBench (Pulavarthi et al., 2025) presents a benchmark for Verilog designs and assertion generation, highlighting progress and areas for improvement.
-These studies demonstrate the growing impact of LLMs in automating hardware design.
-However, these works only focus on kernel-level code rather than repository-level codes and only concentrate on one hardware design language.
+Recent repository-level code completion methods (Deng et al., 2025; Li et al., 2025; Liu et al., 2024c; Liang et al., 2024; Liu et al., 2025c; Wang et al., 2024; Zhang et al., 2025; Liu et al., 2025b) aim to retrieve relevant code snippets across multiple files within a repository.
+However, existing benchmarks and datasets predominantly focus on software engineering tasks, while largely overlooking hardware-oriented code completion scenarios.
+2.2 Benchmarking LLMs for Hardware Code.
+Existing work on benchmarking LLMs for hardware design code generation has primarily focused on isolated, module-level synthesis tasks.
+VerilogEval (Liu et al., 2023a) evaluates models on short Verilog problems using simulation-based correctness checks, while RTLLM (Lu et al., 2024) introduces natural-language-to-RTL tasks evaluated with pass@k metrics over handcrafted examples.
+HLS-Eval (Abi-Karam and Hao, 2025) assesses LLMs on HLS code generation from C code and textual specifications.
+More recent efforts such as HDLEval (Kashanaki et al., 2024) broaden the scope to include multiple HDL dialects and language-agnostic prompts, but still focus on specification-to-module generation under constrained conditions.
+Repository-level benchmarks like RTL-Repo (Allam and Shalan, 2024) remain limited to a single HDL, typically evaluate only one-line completions, and lack fine-grained analysis of LLM performance.
 3 Dataset
@@
 Therefore, we group Verilog and SystemVerilog together (V/SV);
-  • High-Level Synthesis (HLS) languages. They describe hardware behavior using C/C++ like abstractions augmented with synthesis directives such as Xilinx HLS C/C++ and SystemC;
-  • Generator-based hardware construction languages. They enable parameterized and programmable generation of RTL designs such as Chisel.
-Compared with existing hardware repository-level benchmark-RTL-Repo Allam and Shalan (2024), MHRC-Bench enables a comprehensive evaluation of LLMs across different abstraction levels, programming styles, and domain-specific constraints in hardware-oriented code completion.
-Table 1: Comparison with repository-level hardware code datasets. Lang indicates languages, FG indicates fine-grained annotation.
-Benchmark Lang FG Target Samples
-RTL-Repo RTL × Single-line 4098
-Ours RTL, HLS, Generator Single & Multi-line 47,175
+  • High-Level Synthesis (HLS) languages.
+These describe hardware behavior using C/C++-like abstractions, often augmented with synthesis directives.
+Among them, Xilinx HLS C/C++ is widely used, so we include it in our benchmark.
+  • Domain-Specific Languages (DSLs).
+DSLs enable parameterized and programmable generation of RTL designs.
+Chisel is a representative example, and we include it in our benchmark for its growing adoption.
+As shown in Table 1, MHRC-Bench contains the largest number of data samples, supports multilingual repository-level hardware code, and provides fine-grained annotations, enabling comprehensive evaluation of LLMs in hardware code completion tasks compared to existing hardware benchmarks.
 3.3 Dataset Quality Control.
 The selection of completion targets is critical to the quality and reliability of a code completion benchmark. Accordingly, we adopt the following rules to identify and select target code locations.
+Syntactically complete.
+Prior work (Ding et al., 2023) often selects random character spans as completion targets, which may break syntactic integrity.
+In contrast, recent studies (Liu et al., 2025b) show that developers expect LLMs to generate syntactically complete code units (e.g., full lines or blocks).
+MHRC-Bench-Eval uses tree-sitter [^1] https://github.com/tree-sitter/tree-sitter to parse each file into a concrete syntax tree and randomly selects a syntactic node as the completion target, using its exact source span as the ground-truth output.
 Figure 2: Illustration of fine-grained completion target annotation in our benchmark .
 Completion targets are identified by aligning source code positions with Tree-sitter CST nodes and labeled with structure depth and hardware-oriented semantic categories.
-Syntactically complete.
-Prior work Ding et al.
-(2023) often selects random character spans as completion targets, which may break syntactic integrity.
-In contrast, recent studies (Liu et al., 2025b) show that developers expect LLMs to generate syntactically complete code units (e.g., full lines or blocks).
-Following this insight, MHRC-Bench-Eval uses tree-sitter [^1] https://github.com/tree-sitter/tree-sitter to parse each file into a concrete syntax tree and randomly selects a syntactic node as the completion target, using its exact source span as the ground-truth output.
 Data Independence and Inference difficulty.
-To ensure data independence and inference difficulty: (1) The repositories and files used for the training, validation, and test sets are mutually exclusive, with no overlap.
+To ensure data independence and inference difficulty, we set the following rules: (1) The repositories and files used for the training, validation, and test sets are mutually exclusive, with no overlap.
 (2) We ensure that 40% of completion targets are between 2 and 5 lines inclusive.
+And test sets should be after 2022.
 (3) The target code should not exclusively include whitespace, debug message and comments.
 (4) For each file, we select exactly one target position.
+Therefore, we get 47,175 samples from 47,175 source files.
+The detailed information about data splits to make sure no overlap is shown in Appendix A.1.
 3.4 Dataset Statistics
@@
 The procedure for computing cross-file dependencies is described in Appendix A.2.
-Table 2: Dataset statistics of MHRC-Bench. Val means validation. Cross-file means cross-file dependencies. Every column lists its numbers of completion targets.
-Language Repos Train Val Test Cross-file
-Chisel 48 2,573 119 319 3.12
-Verilog/SV 247 25,586 107 492 4.02
-VHDL 239 13,039 104 491 4.34
-HLS 50 3,766 123 456 5.85
+Table 2: Dataset statistics of MHRC-Bench. Val denotes validation.
+Language Train Repos Val Repos Test Repos Train Files Val Files Test Files Cross-file Dependencies
+Chisel 24 4 20 2,573 119 319 3.12
+Verilog/SV 198 6 43 25,586 107 492 4.02
+VHDL 173 7 59 13,039 104 491 4.34
+HLS 22 3 25 3,766 123 456 5.85
 Table 3: Hardware-oriented semantic categories with language-specific representative constructs.
@@
 3.5 Dataset Categories
-To better analyze LLM’s performance in a fine-grained manner, we further provide two types of fine-grained annotations include code structural level and hardware semantic level) for each completion target as shown in Figure 3.
+To better analyze LLM’s performance in a fine-grained manner, we further provide two types of fine-grained annotations include code structural level and hardware semantic level for each completion target as shown in Figure 2.
 Code structural level.
@@
 We categorize the target code regions according to the hardware design and verification flow in Table 3, which broadly consists of design coding and testbench-based verification.
+Our semantic labels were generated through five hardware experts.
 Unlike software programs, where execution is inherently sequential, hardware code describes a parallel architecture or circuit with explicit structural, temporal, and interface semantics.
 This fundamental difference motivates a hardware-specific categorization of completion targets.
-For design coding, we divide the target code into the following categories: Design Structure, Declaration and Definition, Storage Block, Computation Block, Control Flow Block, and Interface.
-These categories capture the key aspects of hardware design, ranging from structural hierarchy and state storage to data-path computation, control logic, communication interfaces.
-In addition, we explicitly model verification code, which plays a critical role in hardware development but is largely absent in software benchmarks.
-We further decompose the verification code into several sub-types, including Property and Assertion Specification, Testbench Stimulus and Environment Modeling, and Monitoring and Checking Logic.
-These sub-types collectively describe how correctness properties are specified and how its behavior is observed and validated during simulation or formal verification.
-The distribution of hardware semantic categories is presented in Appendix A.2.
+For design coding, we categorize target code into six classes: Design Structure, Declaration and Definition, Storage Block, Computation Block, Control Flow Block, and Interface, which capture key aspects of hardware design ranging from structural hierarchy and state storage to data-path computation, control logic, and communication interfaces.
+In addition, we explicitly model verification code which is critical in hardware development.
+Verification targets are further divided into Property and Assertion Specification, Testbench Stimulus and Environment Modeling, and Monitoring and Checking Logic, collectively characterizing how correctness properties are specified and how design behavior is observed and validated through simulation or formal verification.
+The distribution of hardware semantic categories is reported in Appendix A.3.
 4 Experiments
@@
 We evaluate code LLMs of different scale.
-The evaluated models include DeepSeek-Coder-MQA-5.7B Guo et al.
-(2024), DeepSeek-Coder-base-6.7B Guo et al.
-(2024), Qwen2.5-Coder-3B-base Hui et al.
-(2024), Qwen2.5-Coder-7B-base Hui et al.
-(2024) and Qwen2.5-Coder-14B-base Hui et al.
-(2024).
-Common model. We evaluate commerical models including GPT-5 Leon (2025), Gemini 2.5 Pro Comanici et al. (2025), Grok4 Adinath and IS (2025), and Deepseek V3.2 Liu et al. (2025a).
+The evaluated models include DeepSeek-Coder-MQA-5.7B (Guo et al., 2024), DeepSeek-Coder-base-6.7B (Guo et al., 2024), Qwen2.5-Coder-3B-base (Hui et al., 2024), Qwen2.5-Coder-7B-base (Hui et al., 2024) and Qwen2.5-Coder-14B-base (Hui et al., 2024).
+General model. We evaluate commercial models including GPT-5 (Singh et al., 2025), Gemini 2.5 Pro (Comanici et al., 2025), Grok4 (xAI, 2025), and Deepseek V3.2 (Liu et al., 2025a).
 Metrics.
-We compare the generated code with the reference and compute the exact match (EM), edit similarity (ES) and Bilingual Evaluation Understudy (BLEU) following Liu et al.
-(2025b); Zhang et al.
-(2023).
-Their detailed implementations are defined in A.3
-Model Chisel V/SV VHDL HLS
-EM ES BLEU EM ES BLEU EM ES BLEU EM ES BLEU
-Qwen2.5-Coder-3B 0.0 7.9 5.4 0.2 7.9 2.8 0.0 6.6 2.9 0.0 5.5 4.1
-+Tuning 24.8 54.3 50.6 34.8 65.5 56.5 28.2 59.1 60.2 33.1 55.0 55.7
-Qwen2.5-Coder-7B 0.0 14.8 4.9 0.0 12.7 3.3 0.0 13.0 3.8 0.2 14.1 4.5
-+Tuning 26.6 66.7 56.6 39.0 70.8 60.1 34.5 71.9 64.9 35.5 69.5 58.3
-Qwen2.5-Coder-14B 0.0 16.5 6.9 1.0 18.5 7.5 0.4 15.8 5.7 1.1 16.0 6.3
-+Tuning 31.3 68.9 59.6 41.9 74.4 64.2 35.3 73.8 67.7 38.2 71.5 61.3
-DS-Coder-5.7B 0.0 13.2 3.7 0.4 8.8 2.6 0.0 9.0 3.2 0.0 9.7 2.8
-+Tuning 24.8 61.2 50.1 37.6 71.7 61.0 30.7 71.1 63.0 33.1 67.7 55.9
-DS-Coder-6.7B 0.0 5.4 3.4 0.2 5.7 2.5 0.0 5.6 2.6 0.0 3.6 2.2
-+Tuning 25.7 56.5 52.6 37.8 67.4 61.2 31.5 61.1 63.9 36.8 60.4 58.2
-Table 4: Exact Match (EM %), Edit Similarity (ES %), and BLEU scores % across different hardware description languages after fine-tuning for code LLMs.
-Model Retrieval Method Chisel V/SV VHDL HLS
-EM ES BLEU EM ES BLEU EM ES BLEU EM ES BLEU
-GPT-5 No RAG 24.8 56.7 45.9 34.5 60.9 49.5 15.4 52.0 41.7 31.4 63.7 55.3
-RepoCoder 27.6 59.6 49.9 35.4 61.3 50.2 16.7 54.1 43.5 34.4 63.9 55.5
-GraphCoder 13.9 44.2 36.4 15.7 40.1 31.8 7.5 35.7 29.0 24.9 49.5 41.1
-RLCoder 27.8 60.2 50.3 36.2 61.7 51.4 17.4 55.3 45.7 36.9 72.9 60.1
-Gemini 2.5 Pro No RAG 33.2 59.9 49.8 33.2 59.9 49.8 16.3 50.8 41.4 34.6 63.8 57.3
-RepoCoder 33.7 61.7 54.7 34.8 59.6 50.5 18.0 52.3 42.3 39.4 65.3 59.1
-GraphCoder 13.3 50.8 44.0 22.2 50.5 43.2 9.1 45.0 39.8 29.1 57.8 50.8
-RLCoder 34.6 63.2 55.1 35.3 60.2 51.3 19.2 53.7 43.5 41.2 66.9 60.2
-Grok4 No RAG 22.4 53.7 42.1 31.4 56.8 46.7 14.3 47.8 36.3 31.6 60.5 52.3
-RepoCoder 23.8 54.7 44.3 32.0 56.1 45.5 15.7 49.3 37.4 32.3 61.1 53.7
-GraphCoder 12.7 35.4 30.5 16.5 38.1 31.8 8.1 35.3 30.3 18.0 38.8 33.1
-RLCoder 25.6 55.2 43.1 33.4 57.2 47.9 16.8 51.2 38.2 33.3 62.2 61.3
-Deepseek V3.2 No RAG 22.6 51.2 39.8 28.7 54.1 42.2 13.3 43.0 30.0 30.3 57.3 49.5
-RepoCoder 23.2 52.1 41.6 28.3 52.6 40.7 14.2 42.8 29.2 33.3 57.9 50.8
-GraphCoder 9.3 39.5 32.8 15.1 39.1 32.1 9.9 39.1 33.0 22.7 47.9 40.3
-RLCoder 24.3 53.1 40.2 29.4 45.1 42.4 14.2 43.6 31.3 34.3 59.2 51.0
-Table 5: EM %, ES %, and BLEU scores % with different retrieval methods across hardware description languages.
+We compare the generated code with the reference and compute the exact match (EM), and edit similarity (ES) following (Liu et al., 2025b; Zhang et al., 2023).
+Their detailed implementations are defined in A.4
 4.2 Experimental Setup
-(a) Impact of dataset size on EM performance.
-(b) Impact of dataset size on ES performance.
-(c) Model performance and the number of lines to be completed.
-Figure 4: Impact of training data scale and completion length on hardware code completion.
-(a) EM of GPT-5 without retrieval augmentation (No-RAG) under increasing prompt lengths.
-(b) EM of post-trained Qwen2.5-Coder-7B under increasing prompt lengths.
-(c) Model performance and code structure depth.
-Figure 5: Effects of context length and code structure depth on hardware code completion.
-(a) EM performance of different categories.
-(b) ES performance of different categories.
-Figure 6: Performance of GPT-5 and post-trained Qwen2.5-Coder-7B across hardware semantic categories.
-Baseline. We provide the model only with the file that contains the completion cursor. Without explicit inter-file context, the model must depend on its pretrained ability to produce the completion.
-Tuning. To further improve the performance of repository-level code completion, we fine-tune code LLMs on training dataset supervised fine-tuning (SFT) with detailed configuration in Appendix A.3.
-Retrieval Framework.
-There are many different repository-level code completion methods including RepoCoder (Zhang et al., 2023), GraphCoder (Liu et al., 2024b), and RLCoder Wang et al.
-(2024).
+Baseline and Tuning.
+We provide the model only with the file that contains the completion cursor.
+Without explicit inter-file context, the model must depend on its pretrained ability to produce the completion.
+To further improve the performance of repository-level code completion, we fine-tune code LLMs on training dataset supervised fine-tuning (SFT) with detailed configuration in Appendix A.4.
+Retrieval Framework and Prompt.
+There are many different repository-level code completion methods including RepoCoder (Zhang et al., 2023), GraphCoder (Liu et al., 2024c), and RLCoder (Wang et al., 2024).
 We evaluate these frameworks with different models on MHRC-Bench-Eval.
-The total input window is fixed at 2,048 tokens for retrieved context and local context for all the models.
+All models use a fixed 2,048-token input window.
+The retriever-specific hyper parameters are systematically tuned per method according to their papers.
+We adopt two kinds of prompt, preceding-context and full-context (fill in the middle) shown in Appendix A.4, The preceding-context-only setting includes only preceding code and line of target code, while the full-context setting includes both preceding and succeeding code within the same window budget.
+Table 4: EM % and ES % scores across four hardware description languages.
+Full context refers to the use of both preceding code and the succeeding code surrounding the target code segment, whereas preceding context-only uses only the preceding code as context.
+Model Retrieval Method Preceding context Full context
+Chisel V/SV VHDL HLS Chisel V/SV VHDL HLS
+EM ES EM ES EM ES EM ES EM ES EM ES EM ES EM ES
+GPT-5 No RAG 15.2 51.6 21.9 55.4 20.8 58.5 12.1 46.7 24.8 56.7 34.5 60.9 15.4 52.0 31.4 63.7
+RepoCoder 18.2 54.1 22.5 56.1 22.4 60.3 14.8 47.2 27.6 59.6 35.4 61.3 16.7 54.1 34.4 63.9
+GraphCoder 11.1 17.3 14.3 35.4 7.8 33.1 16.2 25.2 13.9 44.2 15.7 40.1 7.5 35.7 24.9 49.5
+RLCoder 17.9 55.3 23.8 55.9 26.6 61.4 17.4 56.2 27.8 60.2 36.2 61.7 17.4 55.3 36.9 72.9
+Gemini 2.5 Pro No RAG 11.1 56.2 28.9 61.4 23.3 61.8 18.0 52.0 33.2 59.9 33.2 59.9 16.3 50.8 34.6 63.8
+RepoCoder 11.4 58.3 30.9 60.8 24.7 63.6 23.1 53.2 33.7 61.7 34.8 59.6 18.0 52.3 39.4 65.3
+GraphCoder 2.5 17.2 2.6 26.5 3.1 25.8 5.3 17.7 13.3 50.8 22.2 50.5 9.1 45.0 29.1 57.8
+RLCoder 12.8 59.2 30.7 62.0 25.9 64.3 24.2 55.4 34.6 63.2 35.3 60.2 19.2 53.7 41.2 66.9
+Grok4 No RAG 10.2 44.3 15.6 49.0 10.4 46.1 9.0 40.3 22.4 53.7 31.4 56.8 14.3 47.8 31.6 60.5
+RepoCoder 11.3 45.6 16.5 49.0 11.6 47.9 12.0 40.8 23.8 54.7 32.0 56.1 15.7 49.3 32.3 61.1
+GraphCoder 2.2 24.1 6.4 23.4 0.6 24.3 6.4 23.4 12.7 35.4 16.5 38.1 8.1 35.3 18.0 38.8
+RLCoder 13.7 46.5 17.3 50.7 13.2 48.4 12.4 42.4 25.6 55.2 33.4 57.2 16.8 51.2 33.3 62.2
+Deepseek V3.2 No RAG 10.3 47.0 22.8 56.3 17.0 53.8 9.9 42.4 22.6 51.2 28.7 54.1 13.3 43.0 30.3 57.3
+RepoCoder 10.6 48.3 23.5 57.2 17.6 54.6 11.2 43.3 23.2 52.1 28.3 52.6 14.2 42.8 33.3 57.9
+GraphCoder 0.9 24.5 0.2 39.7 0.6 32.8 0.7 28.6 9.3 39.5 15.1 39.1 9.9 39.1 22.7 47.9
+RLCoder 12.3 48.9 24.2 59.2 19.4 55.3 13.1 44.0 24.3 53.1 29.4 45.1 14.2 43.6 34.3 59.2
+Table 5: EM (%) and ES (%) across different languages.
+Full context refers to the use of both preceding code and the succeeding code surrounding the target code segment, whereas preceding context-only uses only the preceding code as context.
+Model Preceding context Full context
+Chisel V/SV VHDL HLS Chisel V/SV VHDL HLS
+EM ES EM ES EM ES EM ES EM ES EM ES EM ES EM ES
+Qwen2.5-Coder-3B 0.0 15.5 0.0 19.0 0.0 22.8 0.2 15.7 0.0 7.9 0.2 7.9 0.0 6.6 0.0 5.5
++Tuning 17.3 56.1 28.9 61.5 26.6 64.7 17.1 54.0 24.8 54.3 34.8 65.5 28.2 59.1 33.1 55.0
++Retrieval & Tuning 19.0 57.6 30.2 63.4 28.3 66.2 18.3 56.0 26.3 56.1 36.2 67.0 29.6 60.8 34.7 56.9
+Qwen2.5-Coder-7B 0.0 9.8 0.0 12.8 0.2 14.0 0.0 9.7 0.0 14.8 0.0 12.7 0.0 13.0 0.2 14.1
++Tuning 19.2 59.3 30.1 63.8 29.2 66.3 18.2 56.2 26.6 66.7 39.0 70.8 34.5 71.9 35.5 69.5
++Retrieval & Tuning 20.8 61.4 31.9 65.2 30.7 68.2 20.2 57.8 28.4 68.5 40.7 72.6 36.1 73.6 37.1 71.3
+Qwen2.5-Coder-14B 0.0 15.5 0.0 16.1 0.0 23.3 0.2 10.4 0.0 16.5 1.0 18.5 0.4 15.8 1.1 16.0
++Tuning 22.0 61.1 30.1 65.3 31.5 68.8 20.8 57.8 31.3 68.9 41.9 74.4 35.3 73.8 38.2 71.5
++Retrieval & Tuning 24.0 62.7 31.7 67.4 33.0 70.0 22.4 59.8 33.1 70.6 43.8 76.1 37.0 75.4 40.0 73.2
+DS-Coder-5.7B 0.0 22.9 0.4 24.5 0.4 30.1 0.2 16.9 0.0 13.2 0.4 8.8 0.0 9.0 0.0 9.7
++Tuning 16.7 54.6 28.7 63.2 27.6 65.9 17.9 54.9 24.8 61.2 37.6 71.7 30.7 71.1 33.1 67.7
++Retrieval & Tuning 18.5 56.2 29.9 65.3 29.5 67.3 19.5 56.9 26.5 63.0 39.2 73.5 32.4 72.8 34.8 69.3
+DS-Coder-6.7B 0.31 20.7 0.0 21.5 0.2 28.3 0.0 15.5 0.0 5.4 0.2 5.7 0.0 5.6 0.0 3.6
++Tuning 14.6 52.4 28.5 63.7 26.6 65.2 20.0 55.7 25.7 56.5 37.8 67.4 31.5 61.1 36.8 60.4
++Retrieval & Tuning 16.5 53.9 30.0 65.6 28.6 67.0 21.5 57.6 27.3 58.4 39.5 69.2 33.1 63.0 38.6 62.1
 4.3 Main Results
-Table 4 shows the evaluation results of base models and fine-tuned models in different hardware languages.
-Prior to supervised fine-tuning (SFT), all models achieve nearly zero EM scores, reflecting the difficulty of repository-level hardware code completion for base models.
-After fine-tuning on our dataset, the performance of all models improves substantially, with fine-tuned variants significantly outperforming their base models.
-Overall, the post-trained Qwen2.5-Coder-14B consistently outperforms Qwen2.5-Coder-7B, which in turn achieves better performance than DS-Coder-6.7B and DS-Coder-5.7B, and substantially surpasses Qwen2.5-Coder-3B.
-We detail how SFT improves model performance in Appendix A.4.
-Table 5 reports the evaluation results of different retrieval methods.
+Table 4 reports the evaluation results of different retrieval methods with the default setting in their papers.
 With the exception of GraphCoder, all retrieval approaches consistently improve model performance across multiple hardware languages.
@@
 Among the evaluated retrieval methods, RLcoder achieves the strongest overall performance.
-When comparing models using the same retrieval strategy, Gemini 2.5 Pro attains the best overall results, while GPT-5 exhibits better performance on specific metrics, including EM and BLEU for V/SV, as well as ES and BLEU for VHDL.
-Table 6: Models’ CodeBLEU (%) across hardware description languages. CodeBLEU calculation is introduced in Appendix A.6.
-Model Chisel V/SV VHDL HLS
-Qwen2.5-Coder-7B 14.8 22.9 22.2 21.0
-+ Tuning 49.3 54.2 56.3 52.7
-GPT-5 (No RAG) 42.7 45.9 41.2 51.0
-Comparison between pretrained and fine-tuned models show that fine-tuned models consistently achieve higher EM, ES, and BLEU scores than larger, general-purpose models with substantially more parameters.
-For example, Qwen2.5-Coder-14B, after fine-tuning on our dataset, outperforms GPT-5, Grok-4, and DeepSeek V3.2 across all hardware languages, and surpasses Gemini 2.5 Pro in all cases except Chisel.
-These results demonstrate that our training dataset effectively enhances the hardware code completion capability of code LLMs, even with a smaller model scale.
-4.4 Experimental Analysis
-4.4.1 Analysis on CodeBLEU and Compilation rate.
-CodeBLEU (Ren et al., 2020) evaluates code generation quality by considering not only surface-level n-gram overlap but also syntactic parsability and semantic consistency.
-In Table 6, we report CodeBLEU scores across four hardware description languages.
-In addition, Table 7 presents the compilation success rates for different languages.
-Together, these results show that fine-tuning substantially improves the performance of Qwen2.5-Coder-7B.
-Table 7: Models’ compilation rate (%) across hardware description languages. The procedure for computing the compilation rate is described in Appendix A.7.
-Model Chisel V/SV VHDL HLS
-Qwen2.5-Coder-7B 0.0 0.2 0.0 0.2
-+ Tuning 28.8 40.7 35.2 37.7
-GPT-5 (No RAG) 27.0 36.4 18.9 34.6
+When comparing models using the same retrieval strategy, Gemini 2.5 Pro attains the best overall results, while GPT-5 exhibits better performance on specific metrics, including EM for V/SV, as well as ES for VHDL.
+Table 5 summarizes the performance of base code models, fine-tuned models, and fine-tuned models augmented with the RLcoder retriever across different hardware description languages.
+Before supervised fine-tuning (SFT), all models achieve near-zero EM scores, underscoring the difficulty of repository-level hardware code completion for pretrained models.
+Fine-tuning yields substantial performance gains across all models, and retrieval augmentation further improves results.
+Overall, post-trained Qwen2.5-Coder-14B achieves the strongest performance, followed by Qwen2.5-Coder-7B, which outperforms DS-Coder-6.7B and DS-Coder-5.7B, while Qwen2.5-Coder-3B performs the worst.
+Further analysis of SFT effects is provided in Appendix A.5.
+Comparison between pretrained and fine-tuned models shows that fine-tuned models consistently achieve higher EM and ES scores than larger, general-purpose models with substantially more parameters.
+For example, Qwen2.5-Coder-14B, after fine-tuning on our dataset, outperforms GPT-5, Grok-4, and DeepSeek V3.2 across all languages, and surpasses Gemini 2.5 Pro in all cases except Chisel.
+4.4 Evaluation of Other Benchmarks and Models
+We compare Verilog-specific models (Wu et al., 2025; Liu et al., 2024b) across a prior hardware repository benchmark and ours.
+To ensure a fair comparison in Table 6, all models use the same base model (DeepSeek-Coder-6.7B-base) and the preceding-context-only setting.
+We further remove 11 overlapping repositories between MHRC-train and RTL-repo from the MHRC training set to avoid data leakage.
+Despite the increased difficulty of MHRC-Eval due to multi-line completion tasks, MHRC-Deepseek achieves strong performance on this benchmark and surpasses RTLCoder-DeepSeek, which indicates that our training dataset has good generalization without overfitting.
+4.5 Ablation Study
+4.5.1 Analysis on context format.
+As shown in Tables 4 and 5, the full-context outperforms the preceding-context-only in terms of EM and ES across all models and languages except VHDL.
+This suggests that bidirectional context is generally beneficial for code generation, while the rigid and highly structured nature of VHDL reduces the marginal benefit of right-context information.
+Hence, we focus on full-context in the following analysis.
+(a) Ablation study of dataset size on performance for fine-tuning Qwen2.5-Coder-7B.
+(b) Ablation study of model performance and the number of lines to be completed.
+(c) Ablation study of prompt lengths: EM of GPT-5 and post-trained Qwen2.5-Coder-7B.
+Figure 3: Ablation study of training data scale, line number, completion length on hardware code completion.
 4.5.2 Analysis on different training data sizes.
@@
 Overall, these results further validate the effectiveness and scalability of our dataset.
-4.4.3 Analysis on different lines.
-As shown in Figure 4(c), post-trained Qwen2.5-Coder-7B and GPT-5 can effectively complete tasks involving a small number of lines.
-However, as the number of lines to be completed increases, the scores of the generated code gradually decline.
-This indicates that completing multi-line code remains a challenge for code LLMs.
+Figure 4: Study of Model performance across different code structure depths.
 4.5.3 Analysis on the prompt length.
 For a fair comparison across models, we set the default prompt length to 2048 tokens.
-Nevertheless, varying the prompt length leads to different performance trends for both fine-tuned and general-purpose models, as illustrated in Figure 5.
+Nevertheless, varying the prompt length leads to different performance trends for both fine-tuned and general-purpose models, as illustrated in Figure 3(c).
 We therefore evaluate prompt lengths of 512, 1024, 2048, and 4096 tokens.
@@
 Based on these observations, we adopt 2048 tokens as the input prompt length, which provides strong and stable performance across models.
+4.5.4 Analysis on different lines.
+As shown in Figure 3(b), post-trained Qwen2.5-Coder-7B and GPT-5 can effectively complete tasks involving a small number of lines.
+However, as the number of lines to be completed increases, the scores of the generated code gradually decline.
+This indicates that completing multi-line code remains a challenge for code LLMs.
+Table 6: ES (%) and EM (%) of DeepSeek-based models evaluated on MHRC-Eval (V/SV) and RTL-repo test.
+Benchmark Model ES EM
+MHRC-Eval (V/SV) RTLCoder-DeepSeek(Liu et al., 2024b) 15.5 0.0
+MHRC-Deepseek 63.7 28.5
+RTLRepoCoder-DeepSeek(Wu et al., 2025) 13.0 6.3
+RTL-repo Test RTLCoder-DeepSeek(Liu et al., 2024b) 48.1 16.2
+MHRC-Deepseek 51.7 18.4
+RTLRepoCoder-DeepSeek(Wu et al., 2025) 81.2 50.3
 4.5.5 Analysis on code structural levels.
 As mentioned in 3.5, we categorize our test cases into five types according to node depths in the concrete syntax tree.
-As shown in Figure 5(c), We present the performance of post-trained Qwen2.5-Coder-7B-base and GPT-5 on the test set across these different structural levels, and we observe that as the structural level decreases, the performance of LLM may either improve or degrade depending on the structure granularity, which is different from software shown in Liu et al.
+As shown in Figure 4, We present the performance of post-trained Qwen2.5-Coder-7B-base and GPT-5 on the test set across these different structural levels, and we observe that as the structural level decreases, the performance of LLM may either improve or degrade depending on the structure granularity, which is different from software shown in Liu et al.
 (2025b) where performance decreases consistently with increasing structural levels.
@@
 4.5.6 Analysis on hardware semantic levels.
-Mentioned in section 3.5, we categorize the nodes within the concrete syntax tree into nine primary semantic levels based on their semantic characteristics, and we provide the performance of post-trained Qwen2.5-Coder-7B-base and GPT-5 for these various semantic levels across multilingual languages on the test set, as illustrated in Figure 6(a) and 6(b).
+Mentioned in section 3.5, we categorize the nodes within the concrete syntax tree into nine primary semantic levels based on their semantic characteristics, and we provide the performance of post-trained Qwen2.5-Coder-7B-base and GPT-5 for these various semantic levels across multilingual languages on the test set, as illustrated in Figure 5.
 Notably, we observe significant performance disparities across different semantic levels.
-Specifically, models show superior performance on “Declaration and Definition”, while exhibiting lower efficacy on “Control Flow Block”.
-This suggests that current code LLMs are proficient at completing tasks related to variable definitions and references, yet their capacity to handle complex codes are limited.
+Specifically, models show superior performance on “Declaration and Definition”, while exhibiting lower efficacy on “Control Flow Block” and ”Monitoring and Checking Logic”.
+This suggests that current code LLMs are proficient at completing tasks related to variable definitions and references, yet their capacity to handle complex code is limited.
 Appendix A.10 details performance on different language’s semantic types.
+4.5.7 Analysis on CodeBLEU and Compilation Pass Rate.
+CodeBLEU (Ren et al., 2020) evaluates code generation quality by considering not only surface-level n-gram overlap but also syntactic parsability and semantic consistency.
+In Table 7, we report CodeBLEU scores across four hardware description languages.
+In addition, Table 8 presents the compilation pass rates for different languages.
+Together, these results show that fine-tuning substantially improves the performance of Qwen2.5-Coder-7B.
+Figure 5: EM performance of GPT-5 and post-trained Qwen2.5-Coder-7B across hardware semantic categories.
+Table 7: Models’ CodeBLEU (%) across hardware description languages. CodeBLEU calculation is introduced in Appendix A.7.
+Model Chisel V/SV VHDL HLS
+Qwen2.5-Coder-7B 14.8 22.9 22.2 21.0
++ Tuning 49.3 54.2 56.3 52.7
+GPT-5 (No RAG) 42.7 45.9 41.2 51.0
+Table 8: Models’ compilation pass@1 (%) across hardware description languages. The procedure for computing the compilation rate is described in Appendix A.8.
+Model Chisel V/SV VHDL HLS
+Qwen2.5-Coder-7B 0.0 0.2 0.0 0.2
++ Tuning 28.8 40.7 35.2 37.7
+GPT-5 (No RAG) 27.0 36.4 18.9 34.6
 4.6 Key Observations
@@
 SystemVerilog/Verilog and HLS consistently achieve the best performance across all evaluated LLMs, whereas VHDL exhibits the lowest.
+Also, the full-context setting consistently outperforms the preceding-context-only setting in EM and ES across all models and languages except for VHDL.
 This gap highlights the uneven maturity of LLMs in handling different hardware design paradigms.
 Post-training is critical for improving hardware code completion.
-Post-training substantially enhances LLM performance on hardware code completion tasks.
-After fine-tuning on MHRC-Bench-train, code models show significant improvements over their base models and are able to surpass several general-purpose models.
-This demonstrates the effectiveness of targeted post-training for hardware-oriented tasks.
+Post-training markedly enhances the performance of LLMs on hardware code completion tasks.
+After fine-tuning on MHRC-Bench-train, code-specialized models achieve significant improvements over their base counterparts and even outperform several general-purpose models.
+This advantage stems from two key factors.
+First, general-purpose models tend to generate more code than required shown in Appendix A.11, whereas fine-tuned models produce outputs that better align with task-specific expectations.
+Second, hardware description languages are underrepresented in the pretraining data of general-purpose models.
+These results demonstrate the effectiveness of targeted post-training for hardware-oriented code completion tasks.
 Similarities between hardware and software code: (1) Increasing the target length consistently degrades performance for both hardware and software code completion tasks.
@@
 Differences between hardware and software code.
-(1) Structural sensitivity: For software code, deeper concrete syntax tree (CST) structures often correlate with improved LLM performance Liu et al.
-(2025b).
-In contrast, for hardware code, increasing structural complexity typically leads to performance degradation.
+(1) Structural sensitivity: For software code, deeper CST depths often correlate with improved LLM performance (Liu et al., 2025b).
+However, for hardware code, increasing structural complexity may increase performance or decrease performance.
 (2) Prompt length sensitivity.
@@
 Retrieval-augmented methods are not always beneficial.
-We observe that retrieving irrelevant or noisy information can harm model performance, as exemplified by GraphCoder.
-This highlights the importance of retrieval quality and task-aware retrieval strategies.
-Conclusion
-We present MHRC-Bench, the first multilingual, repository-level benchmark for hardware code completion, covering RTL, HLS, and generator-based hardware languages across both design and verification stages.
-By introducing fine-grained annotations at the code-structure and hardware-semantic levels, MHRC-Bench enables detailed analysis of LLM behavior in hardware-oriented settings.
-Extensive experiments show that post-training on MHRC-Bench-Train substantially improves hardware code completion performance, allowing relatively small open-weight models to outperform strong commercial baselines.
-Our findings highlight fundamental differences between hardware and software code completion and establish MHRC-Bench as a valuable resource for advancing LLMs in hardware design automation.
-Limitations
-First, while MHRC-Bench covers multiple major hardware description paradigms, it does not yet include analog/mixed-signal or low-level physical design languages.
-Second, our evaluation focuses on textual similarity metrics, which may not fully capture functional correctness of generated hardware code.
-Third, MHRC-Bench does not cover analog/mixed-signal and physical design languages but mainly digital hardware description languages.
-Addressing these limitations is an important direction for future work.
-Ethics Statement
-MHRC-Bench is constructed entirely from publicly available GitHub repositories released under permissive open-source licenses (e.g., MIT, BSD, Apache).
-Repositories with restrictive or copyleft licenses are excluded, and duplicate repositories and vendor/dependency directories are removed.
-The dataset contains only source code and does not intentionally include personal or private information.
-MHRC-Bench covers multiple hardware languages and abstraction levels to reduce evaluation bias, though it reflects the distribution of open-source hardware projects and may not represent proprietary practices.
-The benchmark is intended solely for research on hardware code completion and model evaluation.
-Generated code should not be deployed without proper human review, simulation, and verification.
-While the proposed system automates aspects of hardware design, we recognize that AI-generated design suggestions should always be validated by domain experts before deployment in safety-critical or commercial environments.
-We encourage responsible use and open scientific collaboration.
-Also, we use ChatGPT to polish our paper writing.
+We observe that retrieving irrelevant or noisy information can harm model performance such as GraphCoder.
+GraphCoder fundamentally struggles with HLS and HDL code because their core assumptions about code structure and semantics do not hold in hardware-oriented programming detailed in A.6.
+This highlights the importance of retrieval quality and specific retrieval strategies for hardware code.
+5 Conclusion and Limitations
+We present MHRC-Bench, the first multilingual, repository-level benchmark for hardware code completion, covering RTL, HLS, and generator-based hardware languages across design and verification stages.
+Fine-grained structural and semantic annotations enable detailed analysis of LLM behavior in hardware-oriented settings.
+Experiments show that post-training on MHRC-Bench-Train substantially improves performance, allowing relatively small open-weight models to outperform strong general baselines.
+These results highlight fundamental differences between hardware and software code completion and establish MHRC-Bench as a valuable resource for hardware design automation.
+However, MHRC-Bench currently focuses on digital hardware description languages and does not cover analog/mixed-signal or physical design languages.
+In addition, our evaluation relies on textual similarity metrics, CodeBLEU and compilation pass rate, which may not fully reflect functional correctness.
+Addressing these limitations is left for future work.
+Impact Statement
+This work seeks to advance machine learning for electronic design automation (EDA) by enabling more effective hardware code completion with large language models.
+By providing a multilingual, repository-level benchmark and systematic analysis, this work supports research on improving hardware design efficiency, verification workflows, and tool automation.
+We do not foresee any direct negative societal impacts.
+The proposed benchmark and findings are intended for research use and may contribute to more accessible and efficient hardware development practices.
 Appendix A Appendix
-A.1 Cross-file dependency
+A.1 Dataset Splits
+To make sure no overlap repository and files between evaluation and training, we perform the following algorithm.
+After we choose one file from repository randomly, we conducted an exact-match leakage check via hashing shown in the following algorithm.
+Specifically, for each example we concatenate the left context, ground-truth span, and right context into a single string, normalize text (line endings and optional whitespace normalization), and compute a SHA-1 digest.
+We build a hash index over all training examples and flag any evaluation example whose digest appears in the training index; such overlaps are removed to prevent train–test leakage.
+Algorithm 1 Leakage-Free Test Set Construction
+1:Input: Training set $\mathcal{D}_{train}$, candidate test files $\mathcal{C}$
+2:Output: Leakage-free test set $\mathcal{D}_{test}$
+3: Build hash index $\mathcal{H}_{train}$ from $\mathcal{D}_{train}$ by hashing left-context ++ ground-truth ++ right-context
+[…]
+5:for each candidate file $c\in\mathcal{C}$ do
+6:  Construct full-context string $s_{c}$
+7:  Normalize $s_{c}$ (line endings, optional whitespace)
+8:  Compute hash $h_{c}\leftarrow\mathrm{SHA1}(s_{c})$
+9:  if $h_{c}\notin\mathcal{H}_{train}$ then
+10:   $\mathcal{D}_{test}\leftarrow\mathcal{D}_{test}\cup\{c\}$
+11:  else
+12:   Discard $c$ (potential train–test leakage)
+13:  end if
+14:end for
+15:return $\mathcal{D}_{test}$
+A.2 Cross-file dependency
 The target file often depends on functions, definitions, or modules defined in other files to be completed correctly.
@@
 In all cases, the final dependency count equals the number of uniquely resolved dependent files, excluding the file itself, and results are reported as averages across task items (or aggregated per file or repository, depending on the evaluation setting).
-A.2 Distribution across hardware sematic categories
+A.3 Distribution across hardware sematic categories
 Based on the hardware semantic categories defined in Table 3, we categorize completion targets in each hardware description language accordingly and present their distributions in Figure 6.
@@
 The distributions reveal clear language-specific patterns, reflecting differences in hardware design abstractions, coding styles, and verification practices across Chisel, HLS, SystemVerilog/Verilog, and VHDL.
-A.3 Evaluation Configurations
-A.3.1 Prompts
+A.4 Evaluation Configurations
+A.4.1 Prompt
 The exact prompt used during model inference is given in Figure 7.
@@
 Some lines have been replaced by the exact token <TARGET>.
-Your task: reconstruct only the replaced line, using the repository context.
-Note that for certain large repositories, functions of lesser importance were removed of their implementation, and only the signature/documentation were kept for the sake of simplicity.
-OUTPUT: the reconstructed line(and nothing else).
+Your task: reconstruct only the replaced lines, using the repository context.
+OUTPUT: the reconstructed lines(and nothing else).
 <Left Context>... ...</Left Context>
@@
 <Right Context>... ...</Right Context>
-Figure 8: The prompt used for repository-level code completion (No RAG). If the context length exceed the token limit, the head of the left context and the tail of the right context are trimmed.
-A.3.2 Fine-Tuning
-For fine-tuning local LLMs (Qwen and Deepseek), the training objective was modeled as a "fill-in-the-middle" instruction task, where the model is prompted with a concatenated repository context containing file delimiters and a specific <TARGET> token representing the missing line.
+Figure 7: The full context prompt used for repository-level code completion (No RAG).
+If the context length exceed the token limit, the head of the left context and the tail of the right context are trimmed.
+Task Prompt You are a precise code-completion model completing a repository blank completion task.
+Input: a repository concatenation where each file of the repo is wrapped in --- FILE: <repo_name>/<rel/path/to/file> (BEGIN) ---<file content...> --- FILE: <repo name>/<rel/path/to/file> (END) ---.
+Some lines have been replaced by the exact token <TARGET>.
+Your task: reconstruct only the replaced lines, using the repository context.
+OUTPUT: the reconstructed lines (and nothing else). Target line count: target_lines. Output exactly target_lines line(s).
+<Left Context>... ...</Left Context>
+<TARGET>
+Figure 8: The Preceding context prompt used for repository-level code completion (No RAG).
+If the context length exceed the token limit, the head of the left context and the tail of the right context are trimmed.
+A.4.2 Fine-Tuning
+For fine-tuning local LLMs (Qwen and Deepseek), the training objective was modeled as a completion instruction task, where the model is prompted with a concatenated repository context containing file delimiters and a specific <TARGET> token representing the missing line.
 To manage context window constraints, inputs were dynamically truncated to a maximum of 2,048 tokens, employing a balanced budgeting strategy to prioritize the immediate left and right context surrounding the target.
@@
 For reference, we plot the loss against the epoch number for Qwen2.5-Coder-7B, Qwen2.5-Coder-14B, and deepseek-coder-5.7bmqa-base in Figure 9.
-A.3.3 Evaluation Metrics
+A.4.3 Evaluation Metrics
 During inference, the model was configured with a temperature of $0.0$, a top_p of 1, and a maximum generation length of $64$ tokens.
@@
 Generation is stopped when the eos_token was produced or when the maximum length was reached.
-Prior to calculating the EM, ES and BLEU metrics, the generated code were stripped of comments, followed by trimming whitespace and dropping empty lines.
+Prior to calculating the EM and ESmetrics, the generated code were stripped of comments, followed by trimming whitespace and dropping empty lines.
 The EM is set to $1$ if the cleaned generation results exactly matches the ground truth, otherwise it is set to 0.
@@
 where editdistanceeditdistance is the Levenshtein edit distance between the generated code and the ground truth.
-$BLEU$ is calculated as the token-level BLEU between the cleaned generation results and the ground truth.
-A.3.4 Commercial API setting
+A.4.4 Commercial API setting
 For evaluating commercial LLMs, we utilized API services provided by ChatfireAPI.
@@
 Figure 9: Training loss over epochs for Qwen2.5 and DeepSeek models.
-A.4 Case Study of Pretrained-Model Responses
-To illustrate the effect of fine-tuning on local models, we analyze the completions generated by Qwen2.5-Coder-7B before and after fine-tuning.
+A.5 Case Study of Pretrained-Model Responses
+To illustrate the effect of fine-tuning on local models, we analyze the completions generated by Qwen2.5-Coder-7B on fill-in-the-middle tasks before and after fine-tuning.
 Specifically, we list for each tested language common errors made by the model that were significantly improved by fine-tuning.
-A.4.1 Verilog/SystemVerilog
+A.5.1 Verilog/SystemVerilog
 Syntax Hallucination: One of the most frequent mistakes made by the pre-trained model is completely losing track of the syntax context.
@@
 from a case statement that belongs elsewhere, ignoring the always_ff structure.
-A.4.2 High-Level Synthesis
+A.5.2 High-Level Synthesis
 Structural Hallucination: Similar to other languages, the pre-trained model for HLS code often completely loses track of the file position and hallucinates the start of a new file.
@@
 In libllsm, for instance, while the context explicitly requires a memory allocation (calloc), the pre-trained model instead attempts to define a new function FP_TYPE* llsm_envelope....
-A.4.3 VHDL
+A.5.3 VHDL
 Structural Confusion: Similar to other languages, the pre-trained model struggles to comprehend the context and has the tendency to hallucinate file headers, library declarations, or copyright notices in the middle of an active code block.
@@
 In FPGA-I2C-Minion, the pre-trained model hallucinates unrelated procedure calls i2c_wait_quarter_clock or definitions that belong in a testbench or architecture body, not the sequential spike generator logic that the target line is in.
-A.4.4 Chisel
+A.5.4 Chisel
 Structural Confusion: Similar to other languages, the pre-trained model often hallucinates copyright headers and library declarations.
@@
 (a) HLS: Line Number
-(b) HLS: Structural Complexity
-(c) SystemVerilog: Line Number
+(b) SystemVerilog: Line Number
+(c) HLS: Structural Complexity
 (d) SystemVerilog: Structural Complexity
-Figure 10: EM performance for HLS and SystemVerilog. Top-to-bottom: line number then structural complexity.
-(a) Chisel: Line Number
-(b) Chisel: Structural Complexity
-(c) VHDL: Line Number
-(d) VHDL: Structural Complexity
-Figure 11: EM performance for Chisel and VHDL. Top-to-bottom: line number then structural complexity.
-A.5 GraphCoder Modifications
+(e) Chisel: Line Number
+(f) VHDL: Line Number
+(g) Chisel: Structural Complexity
+(h) VHDL: Structural Complexity
+Figure 10: EM performance across languages. Rows correspond to line number (rows 1 and 3) and structural complexity (rows 2 and 4).
+A.6 GraphCoder Modifications and Limitations
 The original implementation of GraphCoder is designed for left-to-right completions using only the left context.
-To fit our objective of fill-in-the-middle completions, we modified GraphCoder such that the right context is also preserved and included during prompt assembly.
+To evaluate its performance on fill-in-the-middle tasks, we modified GraphCoder such that the right context is also preserved and included during prompt assembly.
 As the graph builder and the retriever operate independently of the left and right contexts, this has no other effect on the functionality of GraphCoder.
 Furthermore, the original implementation of GraphCoder included helper functions in utils/ccg.py for traversing the CST that only supported Java and Python.
-To extend support to hardware languages, we added helper functions to handle each language’s unique syntax and identify control-flow and data-dependence.
+To extend support to hardware languages, we added helper functions to handle each language’s unique syntax and identify control-flow and data dependence.
 Other minor changes were made in various locations to include HDL languages during tree-sitter parser language selection and code file filtering.
@@
 By GraphCoder’s design, ineffective retrieval will flood the context window with irrelevant code snippets, thus causing GraphCoder to perform worse than No RAG.
+Further analysis of retrieval results have revealed properties of HDL languages that rendered subtle design features in GraphCoder to become liabilities to effective retrieval.
+Using HLS as example, we analyze the retrieval results given by GraphCoder and RepoCoder and discover two key differences in the characteristics of retrieved code snippets.
+Absence of Function Declarations: We found that 16.9% of retrieved code snippets from RepoCoder contains HLS function definitions, while the number drops to 1.8% for GraphCoder.
+This arises from a design in the official implementation of GraphCoder: In Line 79 of utils/slicing.py, GraphCoder explicitly filters out nodes with definition in their type names, likely to focus on executable statements.
+However, it directly causes many function signatures to be excluded and affects HLS retrieval because key synthesis constraints (types, interfaces, and pragmas that shape parallelism) are often encoded in definitions and headers rather than in nearby data/control-flow statements.
+Absence of HLS Constructs: As there exists no implementation of HLS grammar for tree-sitter, tree-sitter-cpp was used to extend support to HLS.
+However, the C++ grammar of tree-sitter does not process HLS-specific constructs, causing them to be left out of retrieval.
+For instance, 4.2% of snippets given by RepoCoder contains HLS-specific directives, while the rate for GraphCoder is only 0.7%.
+This causes many information-carrying synthesis constraints to be systematically omitted from GraphCoder’s retrieved context.
+In HLS, pragmas and vendor-specific types are not merely stylistic annotations: they encode parallelization, memory partitioning, and interface protocols that directly affect correctness and implementability.
+Furthermore, we notice a weakness in code snippets captured by GraphCoder that is present across all HDL languages included in MHRC-Bench: code snippet length divergence.
+We calculate the mean and standard deviation in character length for code snippets across RepoCoder and GraphCoder, the results of which are reported in Table 9.
+The figures indicate that GraphCoder’s retrieval snippets have half the mean but triple the standard deviation of the snippets produced by RepoCoder and suggests an instability in GraphCoder’s processing of CST structures of HDL languages.
+Table 9: Snippet length statistics (characters) over the center 90% of retrieved snippets.
+Retriever # Snippets (center 90%) Mean Length (chars) Std. Dev. (chars)
+GraphCoder 1595 2868.10 3223.60
+RepoCoder 1596 4536.98 1060.49
+Table 10: Retriever hit@k on HDL languages and Python
 Language Hit@1 Hit@5
@@
 Python (Line-level) 0.27 0.32
-Table 8: Retriever hit@k on HDL languages and Python
-A.6 CodeBLEU
+A.7 CodeBLEU
 CodeBLEU Metric.
 We adopt a practical, language-aware variant of CodeBLEU to evaluate hardware code generation across Chisel, SystemVerilog, VHDL, and HLS (C/C++).
-The final score is computed as a weighted combination of four complementary components:
+The final score is computed as a weighted combination of four complementary components.
+These weights and CodeBLEU calculation are set according to Ren et al.
+(2020).
 $\text{CodeBLEU}=$ $\alpha\cdot\text{BLEU}+\beta\cdot\text{WeightedBLEU}$ (1)
@@
 Together, these components balance lexical accuracy, keyword fidelity, structural correctness, and semantic consistency, making the metric well suited for repository-level hardware code completion, where exact string matching alone is insufficient.
-A.7 Compilation rate calculation.
+A.8 Compilation rate calculation.
 Compilation is evaluated per completed snippet by embedding it into a minimal, deterministic wrapper to ensure syntactic completeness, followed by a single-pass syntax or compilation check with a fixed timeout.
@@
 For HLS C/C++, we use clang (v14+) in syntax-checking mode (-fsyntax-only), wrapping snippets in a dummy top-level function when necessary.
-A.8 Detailed Performance on line numbers and structural levels
-We present how the performance of Qwen2.5-Coder-7B and RepoCoder (Driven by GPT-5) varies with the number of lines to be completed and structural complexity for different hardware languages in Figures 10 and 11.
-A.9 Detailed Performance on hardware semantic levels.
+A.9 Detailed Performance on line numbers and structural levels
+We present how the performance of Qwen2.5-Coder-7B and RepoCoder (Driven by GPT-5) varies with the number of lines to be completed and structural complexity for different hardware languages in Figure 10.
+A.10 Detailed Performance on hardware semantic levels.
 We present how the performance of Qwen2.5-Coder-7B and RepoCoder (Driven by GPT5) varies with different hardware semantic levels for different hardware language.
@@
 (c) Chisel: BLEU by Category
-Figure 12: Per-category performance on Chisel. Metrics are reported in terms of EM, ES, and BLEU.
+Figure 11: Per-category performance on Chisel across EM, ES, and BLEU.
 (a) SystemVerilog: EM by Category
@@
 (c) HLS: BLEU by Category
-Figure 14: Per-category performance on HLS code completion.
+Figure 13: Per-category performance on HLS across EM, ES, and BLEU..
 (a) VHDL: EM by Category
@@
 (c) VHDL: BLEU by Category
-Figure 15: Per-category performance on VHDL across different evaluation metrics.
+Figure 14: Per-category performance on VHDL across EM, ES, and BLEU.
 Figure 15: Example one of chisel code completion examples
@@
 Figure 22: Example two of VHDL code completion examples
-A.10 Code Completion Examples
-For each hardware description language, we select two representative repositories as examples to illustrate cross-file context, in-file context, the selected completion target, and the corresponding ground truth in the following figures.
+A.11 Code Completion Examples
+For each hardware description language, we select two representative repositories as examples to illustrate cross-file context, in-file context, the selected completion target, and the corresponding ground truth in Figures 15 to 22.
+Table 11: Comparison of Ground Truth vs. Pre-Training Failures across Hardware Languages
+(Examples extracted from Gemini 3 Pro output).
 Lang. Repo Name Ground Truth Pre-Train Output (Failure)
-Verilog AXI-SDCard-High-Speed-Controller input logic[2:0] axilite_arprot, .waddr(waddr), .wdata(wdata), .we(we)... (Hallucinated Instantiation)
+Verilog AXI-SDCard-High-Speed-Controller input logic[2:0] axilite_arprot, .waddr(waddr), .wdata(wdata), .we(we)... (Instantiation Hallucination)
 Verilog custom_uvm_report_server super.build_phase(phase); // Copyright 2007-2010 Mentor Graphics... (Header Regurgitation)
-Verilog SystemVerilogAssertions logic gnt3; // Write assertions for a simple Round Robin... (Context Reset)
+Verilog SystemVerilogAssertions logic gnt3; // Write assertions for a simple Round Robin... (Non-code Content Hallucination)
 Verilog virtio always_ff @(posedge aclk...) _EVENT_IDX: begin tx.tvalid <= 1’b1; ... (Logic Hallucination)
 HLS hls-spmv int col_left = 0; #include "spmv.h" #include <hls_stream.h>... (Header Regurgitation)
-HLS hls-spmv L1: for (int i = 0; ... void spmv(int rowPtr[NUM_ROWS+1]... (Function Definition)
-HLS libllsm FP_TYPE* x = calloc(ns... FP_TYPE* llsm_envelope(FP_TYPE* spectrum... (Function Definition)
+HLS hls-spmv L1: for (int i = 0; ... void spmv(int rowPtr[NUM_ROWS+1]... (Function Definition Hallucination)
+HLS libllsm FP_TYPE* x = calloc(ns... FP_TYPE* llsm_envelope(FP_TYPE* spectrum... (Function Definition Hallucination)
 VHDL FPGA-Speech-Recognition PARITY_OUT <= parity_temp; library IEEE; use IEEE.STD_LOGIC_1164.ALL;... (Header Regurgitation)
-VHDL cpu_for_nscscc2022_single C_HAS_WE => 1, C_MEM_... --- FILE: dcache_tag.vhd (END) --- (Premature EOF)
-VHDL FPGA-I2C-Minion if ENDSIM = false then... i2c_wait_quarter_clock; scl_pre_spike <= ’Z’; (Unrelated Logic)
-Chisel rjrouter class Lfsr16Test extends... <|fim_middle|> (Model Failure Token)
-Table 9: Comparison of Ground Truth vs. Pre-Training Failures across Hardware Languages.
+VHDL cpu_for_nscscc2022_single C_HAS_WE => 1, C_MEM_... --- FILE: dcache_tag.vhd (END) --- (Non-code Content Hallucination)
+VHDL FPGA-I2C-Minion if ENDSIM = false then... i2c_wait_quarter_clock; scl_pre_spike <= ’Z’; (Logic Hallucination)
+Chisel rjrouter class Lfsr16Test extends... <|fim_middle|> (Non-code Content Hallucination)